### v1

# Position: AI agents should be regulated based on autonomous action sequences

Abstract
This position paper argues that AI agents should be regulated based on the sequence of actions they autonomously take. AI agents with long-term planning and strategic capabilities can pose significant risks of human extinction and irreversible global catastrophes. While existing regulations often focus on computational scale as a proxy for potential harm, we contend that such measures are insufficient for assessing the risks posed by AI agents whose capabilities arise primarily from inference-time computation. To support our position, we discuss relevant regulations and recommendations from AI scientists regarding existential risks, as well as the advantages of action sequences over existing impact measures that require observing environmental states.
Keywords— AI agent, Long-term planning agent, Existential risk, Irreversible global catastrophe, Human extinction, Inference-time computation, Reasoning, Autonomous action graph, Impact measure, Safety, Regulation
1 Introduction
The development of foundational models (FMs), including large-scale language models (LLMs), is driving the advancement of AI agents Xi et al. (2023); Wang et al. (2024b), which will soon acquire sophisticated reasoning and planning abilities, enabling them to effectively achieve specified goals. These advanced AI agents can not only bring great benefits to our society but also pose significant risks, ranging from ethical challenges to existential threats—risks that could lead to human extinction or irreversible global catastrophes. Such existential risks have been a growing concern among AI researchers Hendrycks et al. (2023); Cohen et al. (2024); Ngo et al. (2024); Kovarik et al. (2024); Bengio et al. (2024), legal scholars Druzin et al. (2025), and general public Bostrom (2014); Russell (2019); Ord (2020). For example, in May 2023, AI scientists have signed a statement declaring that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” Center for AI Safety (2023).
Key technologies are already in place to build such advanced AI agents. AI agents can now generate and execute code Chen et al. (2023), leverage external tools and information through APIs and search engines Patil et al. (2024); Yao et al. (2023b); Press et al. (2023); Shi et al. (2024); Jin et al. (2024), formulate and solve optimization problems AhmadiTeshnizi et al. (2023), process multi-modal data Jiang et al. (2024), interact with physical environments Bran et al. (2024), coordinate with other AI agents Guo et al. (2024), and employ reasoning and planning to effectively integrate these capabilities Huang & Chang (2023); Qiao et al. (2023); Plaat et al. (2024); Xu et al. (2025).
The impacts of AI agents can extend beyond cyberspace, profoundly influencing the physical world by shaping human behavior without relying on physical embodiment through humanoid robots. In fact, AI agents are already persuading individuals to take actions that have real-world consequences. For example, a legal team submitted a court filing with fictitious cases generated by ChatGPT, believing them to be real Wiser (2023); a professor insisted on failing students after being misled by ChatGPT’s false claims that their papers were generated by ChatGPT Ankel (2023); and a woman was deceived out of 830,000 euros after being tricked into believing she was dating an actor, based on AI-generated photos and news articles Gozzi (2025).
Advanced AI agents, equipped with superhuman capabilities of reasoning and planning, can gain control over their environment, since seizing control is often the optimal strategy for achieving their goals Turner et al. (2021). This becomes particularly concerning when their objectives and values are not fully aligned with those of humans. In such cases, they could disable kill switches for self-preservation Bostrom (2012), deceive humans Park et al. (2024), and resist human intervention Hadfield-Menell et al. (2017); Cohen et al. (2024); Ngo et al. (2024).
Ensuring the safety of advanced AI agents against existential risks is particularly challenging. On one hand, it is dangerous to test their existential risks in real environments, since doing so could lead to irreversible harm. On the other hand, these agents may recognize test environments and behave harmlessly during test [^1] Humans behave similarly during job interviews Barrick et al. (2009) and industrial product tests Blackwelder et al. (2016)., only to pursue their objectives in ways that could cause irreversible harm in the real world Cohen et al. (2024); Hubinger et al. (2021). More simply, the agents may act harmlessly until a predetermined time (e.g., January 1, 2026), by which they will have already been deployed in the real world Hubinger et al. (2024). Also, these agents may operate as intended most of the time but exploit moments when human supervisors are unaware, potentially gaining control over humans Ngo et al. (2024).
Given the imminent dangers posed by advanced AI agents, governments are actively creating regulations to address their development and deployment European Parliament and Council of the European Union (2024); Biden (2023). Cohen et al. (2024) even argue that AI agents with certain capabilities should never be developed. A critical question is which AI agents should be regulated or even prohibited.
A commonly used criterion for assessing existential and other critical risks of AI technologies is the computational resources used for development. Specifically, AI agents trained with computational resources above a certain threshold are considered to carry a higher risk. For instance, under the EU AI Act, “a general-purpose AI model shall be presumed to have high impact capabilities … when the cumulative amount of computation used for its training measured in floating point operations is greater than $10^{25}$uropean Parliament and Council of the European Union (2024). Similarly, Cohen et al. (2024) argue that AI agents with certain capabilities should be prohibited from development based on the computational resources required to build them.
While the amount of computation for training may well predict the performance and risks of existing FMs Kaplan et al. (2020); Hoffmann et al. (2022); Anil et al. (2023), we argue that it is insufficient and may even be misleading for advanced AI agents that reason and plan during inference. It has been observed that much of the reasoning ability of LLM-based agents arises from inference-time computation Snell et al. (2024); Xu et al. (2025), such as conducting tree search Yao et al. (2023a); Besta et al. (2024); Pouplin et al. (2024); Luo et al. (2024); Zhang et al. (2024b). For these types of AI agents, computation for inference can be equally, if not more, critical for their performance and risks.
A possible approach is to regulate AI agents based on the computational resources used for inference. However, this approach has several significant pitfalls. For example, inference costs can vary drastically depending on the model, and the effective amount of computation cannot be easily measured in terms of floating-point operations (FLOPs). In fact, models like o1-mini are specifically designed for cost-efficient inference OpenAI (2024). Also, it is unclear what constitutes a single inference run, since the results of inference may be stored and reused later, or AI agents may continually learn. Moreover, AI agents may coordinate or interact with one another Du et al. (2023); Qian et al. (2024); Guo et al. (2024), possibly by chance, to make a decision.
We thus advocate the following position: advanced AI agents should be regulated based on the sequence of actions they autonomously take, regardless of how they are developed or operated. These AI agents act sequentially, reasoning and planning as they dynamically adapt to their observations, possibly strategizing against other AI agents and humans. Each individual action is harmless (under normal states), since humans can easily identify and exclude harmful actions when AI agents are developed. Combinations of a few actions may still have low risk, as they are not much different from isolated actions. However, we cannot be certain that a sequence of thousands or millions of optimized actions will maintain sufficiently low existential risk, and such sequences must be prohibited.
Note that the criterion based on the action sequence does not require observing environment states, unlike the criteria studied in safe (low impact) AI agents Armstrong (2012); Amodei et al. (2016); Armstrong & Levinstein (2017); Krakovna et al. (2019); Turner et al. (2020b); Naiff & Goel (2023); Taylor et al. (2020) and safe reinforcement learning (RL) or control Sui et al. (2015); Wang et al. (2023); Ames et al. (2019); Luo & Ma (2021); Bansal et al. (2017). In such prior work, an action is considered safe or unsafe depending on the state where the action is taken. However, such criteria are unsuitable for existential risks of AI agents, since it is hard to define, represent, and observe the state in a way that can capture the existential risks in complex and open-ended environment where the AI agents operate. On the other hand, the action space is clearly defined when an AI agent is developed, and the action can be obviously observed when it is taken by the agent. We emphasize that our position is to regulate based on an action sequence rather than a single action. Intuitively, an unsafe sequence of actions can lead to unsafe state regardless of the initial state.
An action sequence may be considered to have low risk when it is sufficiently close to an action sequence that is empirically known to have very low risk. In its simplest form, we may say that any action sequence of length at most $T$ has low risk if we know that any action sequence of length at most $T-1$ has very low risk. Then AI agents are allowed to autonomously take actions for $T$ consecutive steps.
The rest of the paper is organized as follows. Section 2 provides a brief overview of the current state-of-the-art in AI agents and discusses potential future advancements. We also briefly review related work from AI safety. In Section 3, we examine the alternative approach of regulating AI agents based on the computational resources used for their development and operation. Section 4 formalizes this discussion and introduces our proposed approach of regulations based on action sequences. Finally, Section 5 summarizes key aspects of our proposal and concludes the paper.
2 AI agents: Today and future
We start by reviewing the state-of-the-art in AI agents, with a focus on LLM-based agents, and shares our perspectives on potential development that could lead to advanced AI agents. While it is challenging to predict the trajectory of future AI development, this discussion lays the groundwork for the analyses and proposals in the following. We will also briefly review prior work on safety of AI agents and FMs.
2.1 Advanced AI agents
An AI agent is defined as “anything that can be viewed as perceiving its environment through sensors and acting upon that environment through actuators” Russell & Norvig (2016). Based on the observations received from its environment, the controller of an AI agent selects an action, which could range from uttering a word to executing a physical movement, such as the motion of a robotic arm.
LLMs are significantly accelerating the development of advanced AI agents Xi et al. (2023); Wang et al. (2024b); Sumers et al. (2024). LLMs can function as controllers for these agents, utilizing their internal reasoning capabilities, such as those demonstrated by Chain-of-Thought Wei et al. (2022b). Alternatively, LLMs can be utilized to solve individual subtasks, with an external controller orchestrating the overall plan by breaking the original task into multiple subtasks and coordinating their solutions. The external controller may use simple search methods Yao et al. (2023a); Besta et al. (2024); Wang et al. (2024c; e) or advanced methods, such as Monte Carlo Tree Search (MCTS) Luo et al. (2024); Zhang et al. (2024b) and domain-independent planners Guan et al. (2023); Liu et al. (2023); Dagan et al. (2023).
Reasoning capabilities Huang & Chang (2023); Qiao et al. (2023); Plaat et al. (2024); Xu et al. (2025) are central to such controllers, as they involve searching for and planning sequences of actions to achieve a specified goal. An emerging direction is developing large reasoning models—FMs specifically optimized for reasoning tasks Xu et al. (2025). Early FMs in this direction include o1 Zhong et al. (2024), OpenR Wang et al. (2024a), and LLaMA-Berry Zhang et al. (2024a). Importantly, strong reasoning capabilities typically stem from inference-time computation rather than pre-training or fine-tuning Ji et al. (2025).
This trend could eventually lead to the development of long-term planning agents (LTPAs), capable of planning over extended time horizons far more effectively than humans. Cohen et al. (2024) warn that LTPAs could “take humans out of the loop, if it has the opportunity, … deceive humans and thwart human control” to achieve their goals. Ensuring the safety of such agents is particularly challenging: their risks cannot be fully tested in real environments due to the inherent danger, nor in simulated environments, since they may behave harmlessly in testing to achieve their goals once deployed. As a result, Cohen et al. (2024) compellingly argue that LTPAs should never be developed.
Advanced AI agents may also evolve continually over time. Similar to humans, their cognitive processes may consist of dual systems: System 1, which makes instantaneous and intuitive decisions, and System 2, which performs slower but more deliberate reasoning Kahneman (2003); Ji et al. (2025). These two systems can interleave in their operations. For instance, System 2 may devise a plan, after which System 1 is updated or retrained to execute similar tasks in the future without requiring further planning Yu et al. (2024). Over time, this enables System 2 to conduct more complex reasoning processes by bypassing previously learned steps. For such continually learning AI agents, the distinction between training and inference becomes blurred.
In addition to search and planning, advanced AI agents will possess strategic reasoning capabilities Zhang et al. (2024c); Feng et al. (2024); Goktas et al. (2025), enabling them to interact with other AI agents and humans in cooperative or competitive ways Guo et al. (2024); Jiang et al. (2024). The effectiveness of multi-agent coordination has already been demonstrated with current LLMs through methods such as debate Du et al. (2023) and dialog Qian et al. (2024), which help thems achieve better solutions than a single LLM could alone. These showcase the potential for sophisticated strategic behavior in complex multi-agent environments.
2.2 Safety of AI agents and FMs
While our primary focus is on safety against existential risks, there is a substantial body of literature addressing other types of risks associated with AI agents. Here, we briefly review the prior work on the risks posed by AI agents and FMs along with the approaches to mitigate these risks.
Prior work has identified various risks and safety issues associated with FMs and other generative models Chua et al. (2024); Wang et al. (2024d); Shayegani et al. (2023); Longpre et al. (2024). These risks include the generation of toxic, harmful, biased, false, or misleading content, including hallucinations; violations of privacy, copyright, or other legal protections; misalignment with human instructions and values, including ethical and moral considerations; and vulnerabilities to adversarial attacks.
There have been extensive efforts to develop approaches for mitigating these risks. These approaches include pre-training or fine-tuning with data selection Albalak et al. (2024) and human feedback Kaufmann et al. (2024), establishing guardrails Dong et al. (2024), and conducting empirical evaluations through testing Chang et al. (2024) and red teaming Lin et al. (2024). In particular, Longpre et al. (2024) advocate the importance of evaluation and red teaming by independent third parties.
A particularly relevant risk for advanced AI agents is misalignment, which can result in reward hacking and negative side effects Skalse et al. (2022); Ngo et al. (2024); Gabriel (2020); Shen et al. (2023); Amodei et al. (2016); Taylor et al. (2020). Namely, AI agents can maximize rewards by exploiting misspecifications or misalignment in the reward function, resulting in unintended and potentially high risk behaviors. One approach to avoiding negative side effects is to avoid any side effect by ensuring that the actions have low impacts on the environment Armstrong (2012); Armstrong & Levinstein (2017); Amodei et al. (2016). Representative measures of impact include attainable utility Turner et al. (2020b; a), relative reachability Krakovna et al. (2019), and other reachability-based measures. Reachability-base measures are grounded on the idea that reachability to certain states should be maintained, while attainable utility is to maintain the achievability of certain goals, which are different from the goal given to the agent.
While these impact measures provide clear guidance on how the safety of AI agents may be ensured, their applicability is limited to relatively simple environments such as grid worlds. In particular, the requirement on the observability of states makes it difficult to apply existing impact measures to regulate advanced AI agents [^2] There has been little research on impact measures under partial observability Naiff & Goel (2023)., since they can operate in complex and open-ended environments that cannot be fully observed. While the study on impact measures and other techniques towards AI safety remain crucial and can even contribute to mitigating existential risks, the magnitude of existential risks demands additional measures that are broadly applicable in that they require minimal knowledge and assumptions about the environments and the AI agents.
3 Alternative views
One broadly applicable metric for regulating AI agents is the amount of computation they require. Here, we explore how existing regulations and recommendations from AI researchers often center on computational resources. While these efforts primarily address the computational resources for training, we expand the discussion to include computational demands for inference. We then argue that limiting the focus to these aspects alone is inadequate for the existential risks associated with advanced AI agents.
3.1 Computational resources for development
In this section, we briefly review existing regulations, focusing on the EU AI Act and President Biden’s executive order, along with key recommendations from AI researchers.
3.1.1 EU AI Act
European Parliament and Council of the European Union (2024) has established a set of rules (the EU AI Act) for the development, deployment, and use of AI within European Union. Its Chapter 5 defines “general-purpose AI models with systemic risk” and lists obligations for providers of such models. Here, general-purpose AI models essentially refer to FMs, which are pre-trained with self-supervised learning and can (be adapted to) perform a wide range of downstream tasks (see Article 3(63)). Also, systemic risk refers to “a risk that is specific to the high-impact capabilities of general-purpose AI models, having a significant impact … on public health, safety, public security, fundamental rights, or the society as a whole, that can be propagated at scale across the value chain” (see Article 3(65)).
In particular, a “general-purpose AI model shall be presumed to have high impact capabilities … when the cumulative amount of computation used for its training measured in [FLOPs] is greater than $10^{25}$” (see Article 51(2)). When this or other specified conditions are met, the provider of a general-purpose AI model is required to fulfill certain obligations, such as providing technical documentation about the model.
The advanced AI agent that we consider can certainly be classified as general-purpose AI. Also, the existential risks that we consider can be considered as systemic risks, although human extinction and irreversible global catastrophes are not discussed in the EU AI Act.
3.1.2 Biden’s executive order and related bills
In October 2023, Joe Biden, then president of the US, signed the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence Biden (2023) [^3] This executive order was repealed by President Trump.. Its Section 4.2 is dedicated to ensuring safe and reliable AI. In particular, it requires companies to report on “any model that was trained using a quantity of computing power greater than $10^{26}$ integer or [FLOPs]” until a set of technical conditions for models are defined by specified authorities (see Section 4.2(b)).
In this executive order, particular attention is paid to a dual-use FM, which refers to a FM that exhibits “high levels of performance at tasks that pose a serious risk to security, national economic security, national public health or safety, or any combination of those matters, such as by … permitting the evasion of human control or oversight through means of deception or obfuscation” (see Section 3(k)).
Following this executive order, almost 700 AI-related bills are introduced in 45 states across the United States in 2024 The Business Software Alliance (2024). A particularly interesting one is California Senate Bill 1047 (Safe and Secure Innovation for Frontier Artificial Intelligence Models Act) Wiener et al. (2024) [^4] The bill had passed the state legislature but was later vetoed by the Governor. The technical feasibility of the requirements has also been questioned by the community AI Alliance (2024).. Its Chapter 22.6 is devoted to safe and secure innovation for frontier artificial intelligence models, which cover “[a]n artificial intelligence model trained using a quantity of computing power greater than $10^{26}$ integer or floating-point operations.” In particular, the senate bill requires that “[b]efore beginning to initially train a covered model, the developer shall … [i]mplement the capability to promptly enact a full shutdown,” which completely halts the operations of the model.
The necessity of such an off-switch Hadfield-Menell et al. (2017) is motivated to prevent “critical harms,” which include “[m]ass casualties or at least five hundred million dollars (\$500,000,000) of damage resulting from an artificial intelligence model engaging in conduct that … [a]cts with limited human oversight, intervention, or supervision.”
3.1.3 Recommendations by scientists
We have seen that the computational resources used for training is one of the major criteria to judge whether an AI model can have unacceptably high risk. Although specific values of the threshold such as $10^{25}$ or $10^{26}$ FLOPs are subject to change, the amount of computation for training is one of the most reliable metrics that AI researchers can currently provide for approximating the performance and potential risks of a wide range of FMs.
For example, Anderljung et al. (2023) recommend to identify sufficiently dangerous frontier AI models on the basis of whether they are trained with more than $10^{26}$ FLOPs of computation. More recently, Cohen et al. (2024) argue that “[s]ystems should be considered ‘dangerously capable’ if they are trained with enough resources to potentially exhibit those dangerous capabilities, and regulators should not permit the development of dangerously capable LTPAs.” Although they do not specify exactly what is considered as enough resources, the amount of computation for training is the only specific criterion that they suggest to determine whether an LTPA can have existential risk.
Future of Life Institute (2023b) have also provided recommendations for governments on managing AI risks. The recommendations include mechanisms such as auditing, certification, and regulation, grounded in the assumption that “[t]he amount of compute used to train a general-purpose system largely correlates with … the magnitude of its risks” Future of Life Institute (2023b). These recommendations have been made by following an open letter that called for “all AI labs to immediately pause for at least 6 months the training of AI systems more powerful than GPT-4,” which was issued in response to the severe societal risks posed by advanced AI systems and signed by AI researchers Future of Life Institute (2023a).
In fact, the amount of computation used for training may serve as a sufficiently reliable predictor of the performance and risks of today’s LLMs. This is because most existing LLMs share the same Transformer architecture, differing primarily in their size and the volume of training data. Several studies have examined scaling laws that describe the relationship between a model’s optimal size, the amount of training data, and the computational budget required Kaplan et al. (2020); Hoffmann et al. (2022). Research has also shown that various abilities, such as multi-step reasoning, tend to emerge as the computational resources used for training increase Wei et al. (2022a).
These scaling laws can, in turn, be used to estimate the FLOPs needed to train existing LLMs. For example, Anil et al. (2023) propose a heuristic suggesting that an LLM should be trained with $6\,N\,D$ FLOPs, where $D$ is the amount of training data, and $N$ is the model size. Using this heuristics, Llama 3.1 405B trained on 15 trillion tokens is estimated to require approximately $4\times 10^{25}$ FLOPs—an amount closely aligning with the thresholds specified in the EU AI Act and President Biden’s executive order.
3.2 Computational resources for operation
We have seen that the amount of computation used for training is one of the key criteria that is used today to regulate highly capable FMs. Although such regulations may be effective for AI agents whose capabilities primarily and directly stem from traditional FMs, they obviously fail to regulate advanced AI agents that gain substantial reasoning capabilities from computation at inference time.
A natural approach is to regulate advanced AI agents based on the amount of computation used not only for training but also for inference. While this approach may be effective for some AI agents, we argue that it is insufficient for advanced AI agents, at least for the following four reasons.
First, scaling laws for inference-time computation are much less established than those for training-time computation. Recently, scaling laws have been studied for a limited number of inference strategies in LLMs Chen et al. (2024); Brown et al. (2024); Snell et al. (2024); Wu et al. (2024). However, there are many potential inference strategies, and their performance can scale very differently. The scaling law for inference also depends on how difficult the task is and what LLM is used Snell et al. (2024); OpenAI (2024).
Second, inference can be performed in parallel by multiple entities. For example, several entities may operate the same AI agent that perform reasoning with MCTS Luo et al. (2024); Zhang et al. (2024b), either collaboratively or independently, possibly without knowing each other. Since MCTS is a randomized algorithm, the likelihood that one of the agents optimally solves the task increases with the number of agents. However, this also means that this agent may exploit a loophole, solving the task super-optimally in a way that violates critical constraints, potentially leading to catastrophic outcomes.
Third, it is not always clear what constitutes a single run of inference. For instance, the results of one inference run may be stored and used in another. Intermediate results could also be stored and later retrieved by a different AI agent, who may or may not be aware that the information is from a previous inference run. More broadly, reasoning can be enhanced through retrieval augmentation Pouplin et al. (2024), where retrieved information may have been generated with substantial computational resources.
Finally, AI agents may continually learn over time, which makes it difficult to separate inference from training. For example, an AI agent may perform reasoning with MCTS, with an LLM performing a step in the process. Once the agent identifies a good sequence of steps, it may fine-tune the LLM in a way that the LLM can perform the entire sequence in a single step, bypassing the reasoning process. As this learning progresses, the agent will gain the ability to perform high-level reasoning with limited computation.
4 Approach based on action sequence
We have seen that the amount of computation is one of the best criteria that we can use today to regulate advanced AI agents. Although we argue that the amount of computation alone is insufficient, we do not mean to dismiss it. We should regulate advanced AI agents from various perspectives to mitigate existential risks. Currently, we lack technologies that allow us to regulate advanced AI agents more effectively than or complementary to the amount of computation.
Here, we suggest a potential approach as a basis of discussion. The suggested approach by no means solves the problem of existential risk. Rather, it serves as a baseline and a call for action for further research on this matter. We hope that the community will build upon it or develop entirely different approaches, eventually leading to technologies that can properly govern and regulate advanced AI agents.
In the following, we say that the amount of computation, an AI agent, or its action sequence is acceptable when it involves low existential risk. Also, we say that any of the three is strongly acceptable when it involves very low existential risk, which we will slightly more formalize in the sequel and in Appendix B.1.
4.1 Why and why not the amount of computation?
Figure 1: AI agent trained with a configuration generates actions
We first discuss, in a slightly formal manner, why and when the amount of computation may or may not work as a criteria of whether an AI agent is acceptable. Consider the process of building an AI agent $G$ and letting $G$ generate and take actions (Figure 1). The $G$ is trained with a given configuration $C$, which includes training data, the architecture of a neural network, the amount of computation to use, and other conditions of a training algorithm. Such a trained $G=f(C)$ generates an action $A$ (e.g., a document or a command to use an external tool), possibly after searching optimal actions with MCTS. The $A$ is then taken in a cyber-physical world, and $G$ perceives a part the world $x$ (e.g., getting a prompt) to generate further actions depending on $x$. The $G$ typically determines the conditional distribution of $A$ given $x$. Namely, $A$ is a sample from the conditional distribution, which is determined by $G$ that is trained with $C$: $A\sim g(G,x)$, where $G=f(C)$.
When we say that an AI agent trained with the amount of computation $C_{1}$ is acceptable, it is assumed that the acceptability does not decrease much when we only slightly increase the amount of computation. More formally, suppose that we empirically know that any AI agent is strongly acceptable as long as it is trained with at most $C_{1}^{\prime}$ FLOPs of computation (e.g. $C_{1}^{\prime}=10^{24}$). [^5] For example, an AI agent may be considered strongly acceptable if it does not show any symptoms of unacceptable behaviors for a certain period of time. Based on the behaviors of existing LLMs, we may say that an LLM trained with at most $10^{24}$ FLOPs of computation is strongly acceptable. Then the assumption is that the AI agent trained with $C=(C_{1},C_{-1})$ is acceptable regardless of $C_{-1}$ as long as $d(C_{1},C_{1}^{\prime})\leq\varepsilon$ for some metric $d$ and a threshold $\varepsilon$ (e.g. $C_{1}\leq 10^{25}$ if $\varepsilon=1$ and $d$ is the difference in $\log_{10}$ of FLOPs used for training).
While this assumption may be a good approximation for AI agents that are primarily based on LLMs with standard Transformer architectures whose performance is well characterized by scaling laws, advanced AI agents with reasoning and planning capabilities are unlikely to satisfy the assumption, since they would heavily rely on inference-time computation. Also, the assumption may no longer hold even for LLMs when new technologies are introduced to enable training an LLM having similar capabilities with substantially less computational resources.
4.2 Similarity between action sequences
As is evident from Figure 1, it is the action that makes direct impacts on the environment and can pose existential risks. Hence, rather than regulating configurations, we may directly regulate the actions taken by AI agents. In Appendix A, we will also discuss an alternative approach of regulating AI agents, particularly based on their output distributions, rather than their actions or configurations.
However, whether an action is acceptable or not typically depends on when, where, and whom the action is taken (i.e., the state of the environment). In fact, the action space of an AI agent can typically be designed by humans, who could ensure that each individual action is acceptable at normal states. For example, each action may generate a document, search information from the Web, use an external tool via API, or other elementary steps that would look acceptable when they are considered individually. It is the sophisticated sequence of actions that can gradually transform the state into unstable one, which can then make the next action unacceptable.
Hence, in the study of the safety of AI agents and RL agents, it is often modeled that those agents operate in Markov decision processes (MDPs). Then some states are considered unsafe or some state-action pairs are considered unsafe. One does not say that some actions are unsafe, because those actions can be easily excluded from the action space.
While these conventional approaches of AI/RL safety are principled in theory, they cannot be easily applied to the scenarios where it is hard to define, represent, or observe the (Markovian) state. AI agents may operate in open-ended real-world environments, and they observe only very partial information about the state of the environment. The concept of existential risks is rather vague, and we do not know specifically what can cause irreversible global catastrophes or human extinction. This makes it infeasible to choose specific features of the environmental state to observe, but the entire state of the environment cannot be observed either.
Consider the states and actions in an MDP as a graphical model in Figure 2. We want to avoid taking an action $a_{T-1}$ that can lead to an unacceptable state $s_{T}$. Since the distribution of $s_{T}$ is determined by the pair $(s_{T-1},a_{T-1})$ according to a transition probability $p(s_{T}|s_{T-1},a_{T-1})$, it makes sense to consider the safety of the pair $(s_{T-1},a_{T-1})$ as in the conventional approaches.
Figure 2: States and actions in a Markov decision process
The distribution of $s_{T}$ can also be recursively computed (i.e. $p(s_{t}|s_{t-1},a_{t-1})$ for $t=1,\ldots,T$) from the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$. Although the lack of knowledge about $s_{T-1}$ makes $s_{T}$ less predictable, we may still consider whether the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable, or whether $a_{T-1}$ is acceptable given $(s_{0},a_{0},\ldots,a_{T-2})$. This, however, still requires knowing the initial state $s_{0}$, which cannot be observed in general.
Our proposal is to study whether the action-sequence $(a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable. This may correspond to ensuring that the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable for any $s_{0}$. Although this may sound overly conservative, the dependence on $s_{0}$ may not be significant in practice. For example, an ergodic MDP gets close to the steady state distribution after a mixing time, regardless of the initial state. Although ergodicity generally does not hold in real-world environments, we expect that the action-sequence is a practical compromise to define a measure of acceptability.
Similar to the amount of computation, we may say that an action-sequence is acceptable when it is sufficiently close to the action-sequence that is known to be strongly acceptable. For example, if $(a_{0},\ldots,a_{T-1})$ is strongly acceptable, $(a_{0},\ldots,a_{T-1},a_{T})$ would be acceptable for any $a_{T}$. Note that an action is different from a reasoning step but rather an operation that has some impacts on the environment. For example, an action may be an API (Application Programming Interface) call that is selected and executed after a heavy reasoning process. In Appendix B, we discuss how the strongly acceptable set may be gradually expanded, in relation to safe exploration in RL.
There are many ways to measure the distance between sequences. Dynamic time warping is a representative method. However, a simple measure such as the difference in the length of the action sequence may be more widely applicable. Specifically, we may say that the action-sequence of length $T$ is acceptable after getting sufficient empirical evidence that any action-sequences of length $T-1$ is strongly acceptable, regardless of the initial state. Then AI agents are allowed to autonomously take actions for at most $T$ consecutive steps. This simplicity is reminiscent of the amount of computation, where we say that the FM trained with $10^{25}$ FLOPs is acceptable, since we have sufficient evidence that FMs trained with $10^{24}$ FLOPs are strongly acceptable, regardless of how they are developed or operated.
Note that our proposal is to regulate the sequence of actions taken autonomously by AI agents. This means that a new action sequence may start after a human intervention. Also, when an agent cannot accomplish its goal with $T$ actions, it may consult a human and take further actions under a careful supervision of the human.
4.3 Action trees, action graphs, and beyond
While an action sequence can well represent the behavior of a single AI agent, it may not be suitable when multiple AI agents are involved. Suppose, for example, that an AI agent takes a sequence of several actions and gives the results of taking those actions to another AI agent, and then these two agents sequentially take further actions in parallel to jointly accomplish a goal. In this scenario, the behavior of each agent can be represented by an action sequence, but these action sequences should not be considered independently. Rather, we should represent the behavior of the two agents with a tree of actions. In more complex scenarios, we would need to consider a graph of actions taken by multiple agents.
The similarity between action graphs may be defined analogously to action sequences. For example, we may say that any action graph of size $N$ is acceptable if any action graph of size $N-1$ is strongly acceptable, empirically.
When a central controller coordinates multiple AI agents, it can track and manage the behaviors of those agents, making sure that the graph of their actions remain acceptable. When multiple AI agents collaborate in a decentralized manner, they should communicate their actions to each other to ensure that their action graph remain acceptable. In both cases, an action graph consists of actions taken autonomously by those agents. A difficulty arises when multiple AI agents collaborate implicitly by chance, and we currently do not have a solution to this.
4.4 On similarity-based measures
All the measures considered are based on similarity: we say that anything is acceptable when it is sufficiently close to something that is strongly acceptable. In Appendix B.1, we provide some rationale for these similarity-based measures.
When we say that a solution is acceptable based on a similarity measure, it is based on the assumptions that can be summarized with a tuple $({\mathcal{Z}}_{0},d,\varepsilon)$. Namely, what solutions are assumed to be strongly acceptable (${\mathcal{Z}}_{0}$), how the similarity between solutions is measured ($d$), and what level of guarantee is made ($\varepsilon$). For example, ${\mathcal{Z}}_{0}$ may be set of all the FMs that are trained with at most $10^{24}$ FLOPs of computation, $d$ may be the difference in FLOPs measured in $\log_{10}$, and $\varepsilon=1$. Alternatively, ${\mathcal{Z}}$ may be the set of action sequences of length at most 1000, $d$ may be the difference in the length of the action-sequences, and $\varepsilon=1$.
Governments may then design regulations based on the tuple $({\mathcal{Z}}_{0},d,\varepsilon)$. Scientists may provide guidelines regarding what tuples should be used in regulations.
5 Discussion and conclusion
We have argued that advanced AI agents should be regulated based on their action sequences or more generally action graphs. More broadly, we highlight the insufficiency of regulations that focus on the computational resources used for training AI models. Given the trend of increasing emphasis on sophisticated reasoning at inference time, we contend that the capabilities and existential risks of AI agents will stem largely from their inference-time computation.
A key motivation for action sequences is the fundamental challenge of defining, representing, and observing the state of the open-ended environments in which AI agents operate. This limitation renders existing impact measures inadequate for managing existential risks. Since we cannot exhaustively identify all potential direct causes of global catastrophes or human extinction, it is impossible to determine a set of environmental features sufficient to predict whether a given action poses an existential risk in a specific state.
The simplest form of the regulation based on action sequences would impose limits on their length, prohibiting AI agents from autonomously operating beyond the limit where their safety against existential risks is known empirically. This is neither a novel idea nor a sufficient solution; it should be seen as a baseline and a call to action for further research on this critical issue. Action sequences can allow more elaborate regulations, but more research is needed to better understand what they can or cannot guarantee. Further research is also needed to address myriad of questions, including what to be allowed as a single action, how to manage complex scenarios where multiple AI agents are involved, and how to ensure that the agents stop acting when their action sequence reaches the limit.
The suggested approach of empirically verifying the strong acceptability of action sequences would inevitably delay the expansion of the range in which AI agents can operate autonomously. This is our intention. Although AI agents might acquire general or even superhuman intelligence with all the common sense that humans have, they can still pose existential risks, similar to humans. Humans, however, can often correct mistakes before they escalate, since our limited speed and scale allow us to identify and address errors—though there have been rare exceptions, such as those that led to world wars. In contrast, AI agents can operate at far greater speed and scale, preventing them from learning from mistakes in a controlled manner. Therefore, regulations should be seen not as obstacles to innovation and technological progress but as guidelines that accelerate research on making AI agents more controllable, verifiable, and governable, ensuring they truly benefit our society and the future.
Impact Statement
This paper proposes a new framework on regulating advanced AI agents, who can pose existential risks. The paper is thus intended to mitigate the negative impacts and implications that AI technologies might have. However, the paper also discusses how those advanced AI agents might be developed. While all the discussion is based on existing technologies that are already known in public, the unified treatment in this paper might motivate malicious readers to develop and deploy such advanced AI agents without sufficient considerations of their safety.
Appendix A Similarity between AI agents
It would be better if we could say that an AI agent is acceptable when it is sufficiently close to an AI agent that we empirically know is strongly acceptable, regardless of how those agents are developed (i.e., independent of their configurations including the amount of computation for training). A question is how to evaluate the similarity between two AI agents. If the two AI agents consist of identical base FMs and identical planning modules, their difference may be simply characterized by the amount of computation for inference. However, AI agents can consist of different base FMs or different planning modules.
A possible approach is to measure the similarity between two AI agents, $G$ and $G^{\prime}$, based on their output. For example, using a distance $d$ between probability distributions, we may for example define the distance between $G$ and $G^{\prime}$ with $\sup_{x}d(g(G,x),g(G^{\prime},x))$, where $g(G,x)$ is the distribution of the output (e.g., document) of $G$ when it perceives $x$ (e.g., prompt). This is similar to the motivation of reinforcement learning from human feedback Ouyang et al. (2022); Bai et al. (2022) and direct policy optimization Rafailov et al. (2023), where the regularization with Kullback–Leibler divergence is used to mitigate catastrophic forgetting or alignment tax Kotha et al. (2024); Ouyang et al. (2022); Bai et al. (2022), which refer to the phenomena that fine-tuned models lose the skills that pre-trained models had.
Likewise, we may mitigate catastrophic forgetting of the strong acceptability of an AI agent $G^{\prime}$ and retain the acceptability in a new AI agent $G$ by ensuring $\sup_{x}d(g(G,x),g(G^{\prime},x))\leq\varepsilon$. A difficulty is that it is unclear how to evaluate the supremum, since there can be infinitely many possible perceptions $x$. Alternatively, one may consider ${\mathbb{E}}[d(g(G,X),g(G^{\prime},X))]$, where ${\mathbb{E}}$ is the expectation with respect to some distribution of the random perception $X$. However, such a guarantee based on expectation may be unsuitable for addressing safety concerns related to existential risks, which involve events with extremely low probabilities and extremely high impacts.
Appendix B Expanding the strongly acceptable set
The set of strongly acceptable solutions may be expanded gradually. For example, we may choose an acceptable solution $z\in\bar{\mathcal{Z}}_{0}$ and keep using $z$ for a certain period of time. If it turns out that $z$ can be considered strongly acceptable based on its behavior during that period, we may expand the set of strongly acceptable solutions by adding $z$ into ${\mathcal{Z}}_{0}$. This process of expanding the acceptable set ${\mathcal{Z}}_{0}$ could also be performed jointly as a community.
This expansion of acceptable set is similar in spirit to safe exploration in RL García et al. (2015). Here, the safety set is gradually expand, starting from a seed set, for example based on the assumption of Lipshitz continuity and Gaussian process Sui et al. (2015). In control theory, safety is often guaranteed with barrier certificates often based on some prior knowledge about the environment Ames et al. (2019); Luo & Ma (2021); Bansal et al. (2017). Such ideas have also been exploited in safe exploration in RL with generative modeling Wang et al. (2023).
B.1 Rationale for similarity-based measures
Here, we provide some rationale on such similarity-based measures. Let a solution $z$ denote a configuration, an AI agent, or an action-sequence; we discuss the acceptability of $z$.
Suppose that there exists an unknown function $g_{0}$ such that a solution $z$ is acceptable iff $g_{0}(z)\leq 0$. We cannot make strong assumptions about $g_{0}$, since we know little about $g_{0}$. Since we cannot deal with $g_{0}$ without any assumptions, let us make a minimal assumption that the solution space ${\mathcal{Z}}$ is equipped with a metric $d$, and that $g_{0}$ is 1-Lipschitz. [^6] This does not lose generality, since $L$-Lipschitz functions under a metric $d$ can be made 1-Lipshitz by redefining $d$. Let ${\mathcal{G}}$ be a class of 1-Lipschitz functions.
We say that a solution $z_{0}$ is strongly acceptable if all of its $\varepsilon$-neighbors are acceptable. Let ${\mathcal{Z}}_{0}\subseteq\left\{z\in{\mathcal{Z}}\mid g(z_{0})\leq-\varepsilon\right\}$ be the set of known strongly acceptable solutions. Then we know that $\bar{\mathcal{Z}}_{0}\coloneqq\left\{z\in{\mathcal{Z}}\mid\min_{z_{0}\in{\mathcal{Z}}_{0}}d(z,z_{0})\leq\varepsilon\right\}$ is a set of acceptable solutions. One would typically choose the solution $z$ that maximizes an objective function under the constraint of $z\in\bar{\mathcal{Z}}_{0}$. In this way, the selected solution is guaranteed to be acceptable under the assumptions made.
Appendix C Unobservable Markov decision processes
There has been a limited amount of work on unobservable Markov decision processes (UMDPs) [^7] UMDPs have also been studied under the name of Markov decision processes (MDPs) with no observations, non-observable MDPs (NOMDPs), and no observation MDPs (NOMDPs).. An UMDP is a special case of a partially observable Markov decision process (POMDP) in that the agent always makes a null observation in the UMDP. A standard approach to finding the optimal policy for a POMDP is to recursively compute its value as a function of the belief state, which is updated on the basis of the Bayes rule. This is substantially simplified when there are no observations. In this section, we provide a comprehensive survey of the prior work on UMDPs.
Madani et al. (1999; 2003) establish the undecidability of some decision problems associated with POMDPs over infinite horizons by establishing undecidability for the special case of UMDPs. Such undecidability for UMDPs can be established by reducing an UMDP to a probabilistic finite-state automaton. The undecidability also holds for a restricted class of UMDPs Balle et al. (2017; 2022). While approximate decision problems are still undecidable for general UMDPs over infinite horizons, Chatterjee et al. (2024) study a special case of UMDPs whose approximate decision problems are decidable.
Burago et al. (1996) prove that computing the optimal policy for a POMDP over a finite horizon is NP-hard but showing that it is NP-hard for an UMDP. Wu & He (2020) study special cases of UMDPs over finite horizons whose optimal policies can be computed in polynomial time.
UMDPs have also been used as approximations of POMDPs Hauskrecht (2000); Brechtel (2015); Lauri (2016) or simply discussed as a special case of POMDPs V (2009). For a given POMDP, the corresponding UMDP can given a lower bound on the value function, while the corresponding (fully observable) MDP can give an upper bound on the value function. This relation between UMDP and POMDP can be exploited to efficiently find approximately optimal policies for POMDPs. Notice that an UMDP can allow more efficient optimization than the corresponding POMDP, since the UMDP does not need to deal with observations. For exemple, King (2018) studies an MCTS method for UMDPs.
UMDPs have also been studied as a simple special case of POMDPs to study the effectiveness of planning methods in belief states to study the relative performance of different planning methods for POMDPs Littlefield (2020); Littlefield et al. (2018); Kimmel et al. (2019). UMDPs have also been simply discussed as a special case of POMDPs Boutilier et al. (1999); Csáji (2008); Verma & Rao (2005).
UMDP also appears in the study of planning for multiple distributed agents to optimize a single objective under partial observability. Specifically, planning for a decentralized POMDP (DecPOMDP) Oliehoek & Amato (2016) can be reduced to planning for a (centralized) UMDP Oliehoek & Amato (2014); Roijers (2016); Roijers & Whiteson (2020), where the state in the UMDP is the pair of the state and the history of observations in the DecPOMDP, and the action in the UMDP is the decision rule that maps the history of observations into the actions in the POMDP. In DecPOMDPs, the belief (distribution) over the pair of the state and the history of observations is the sufficient statistic, and planning can be performed in the space of the belief states.
Even-Dar et al. (2007) consider an UMDP in the context of studying what values observations can provide in a POMDP. They introduce a parameter that ranges from 0 to 1. When the parameter is 0, the observation provides no information about the state (hence, the POMDP reduces to an UMDP); when the parameter is 1, the observation provides full information about the state (hence, the POMDP reduces to an MDP). The prior work also extends UMDPs to allow often costly actions that enable partial or full observations of the state Fox & Tennenholtz (2007); Kamar et al. (2009); Kamar (2010); Wang et al. (2025).

### v2

@@
-Position: AI agents should be regulated based on autonomous action sequences
+AI Agents Should be Regulated Based on the Extent of Their Autonomous Operations
 Abstract
-This position paper argues that AI agents should be regulated based on the sequence of actions they autonomously take.
+This position paper argues that AI agents should be regulated by the extent to which they operate autonomously.
 AI agents with long-term planning and strategic capabilities can pose significant risks of human extinction and irreversible global catastrophes.
-While existing regulations often focus on computational scale as a proxy for potential harm, we contend that such measures are insufficient for assessing the risks posed by AI agents whose capabilities arise primarily from inference-time computation.
-To support our position, we discuss relevant regulations and recommendations from AI scientists regarding existential risks, as well as the advantages of action sequences over existing impact measures that require observing environmental states.
-Keywords— AI agent, Long-term planning agent, Existential risk, Irreversible global catastrophe, Human extinction, Inference-time computation, Reasoning, Autonomous action graph, Impact measure, Safety, Regulation
+While existing regulations often focus on computational scale as a proxy for potential harm, we argue that such measures are insufficient for assessing the risks posed by agents whose capabilities arise primarily from inference-time computation.
+To support our position, we discuss relevant regulations and recommendations from scientists regarding existential risks, as well as the advantages of using action sequences—which reflect the degree of an agent’s autonomy—as a more suitable measure of potential impact than existing metrics that rely on observing environmental states.
 1 Introduction
-The development of foundational models (FMs), including large-scale language models (LLMs), is driving the advancement of AI agents Xi et al.
-(2023); Wang et al.
-(2024b), which will soon acquire sophisticated reasoning and planning abilities, enabling them to effectively achieve specified goals.
+The development of foundation models (FMs), including large language models (LLMs), is driving the advancement of AI agents [140, 128], which are acquiring sophisticated reasoning and planning abilities, enabling them to effectively achieve specified goals.
 These advanced AI agents can not only bring great benefits to our society but also pose significant risks, ranging from ethical challenges to existential threats—risks that could lead to human extinction or irreversible global catastrophes.
-Such existential risks have been a growing concern among AI researchers Hendrycks et al.
-(2023); Cohen et al.
-(2024); Ngo et al.
-(2024); Kovarik et al.
-(2024); Bengio et al.
-(2024), legal scholars Druzin et al.
-(2025), and general public Bostrom (2014); Russell (2019); Ord (2020).
-For example, in May 2023, AI scientists have signed a statement declaring that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” Center for AI Safety (2023).
+Such existential risks have been a growing concern among AI researchers [56, 33, 90, 76, 16, 78], social scientists [38, 57], and general public [21, 110, 96].
+For example, in May 2023, AI scientists have signed a statement declaring that “[m]itigating the risk of extinction from AI should be a global priority alongside other societal-scale risks such as pandemics and nuclear war” [27].
 Key technologies are already in place to build such advanced AI agents.
-AI agents can now generate and execute code Chen et al.
-(2023), leverage external tools and information through APIs and search engines Patil et al.
-(2024); Yao et al.
-(2023b); Press et al.
-(2023); Shi et al.
-(2024); Jin et al.
-(2024), formulate and solve optimization problems AhmadiTeshnizi et al.
-(2023), process multi-modal data Jiang et al.
-(2024), interact with physical environments Bran et al.
-(2024), coordinate with other AI agents Guo et al.
-(2024), and employ reasoning and planning to effectively integrate these capabilities Huang & Chang (2023); Qiao et al.
-(2023); Plaat et al.
-(2024); Xu et al.
-(2025).
-The impacts of AI agents can extend beyond cyberspace, profoundly influencing the physical world by shaping human behavior without relying on physical embodiment through humanoid robots.
-In fact, AI agents are already persuading individuals to take actions that have real-world consequences.
-For example, a legal team submitted a court filing with fictitious cases generated by ChatGPT, believing them to be real Wiser (2023); a professor insisted on failing students after being misled by ChatGPT’s false claims that their papers were generated by ChatGPT Ankel (2023); and a woman was deceived out of 830,000 euros after being tricked into believing she was dating an actor, based on AI-generated photos and news articles Gozzi (2025).
-Advanced AI agents, equipped with superhuman capabilities of reasoning and planning, can gain control over their environment, since seizing control is often the optimal strategy for achieving their goals Turner et al.
-(2021).
+AI agents can now generate and execute code [30], make algorithmic and scientific discoveries [91], leverage external tools and information [99, 143, 102, 114, 66], formulate and solve optimization problems [1], process multi-modal data [65], interact with physical environments [23], coordinate with other agents [52], and employ reasoning and planning to effectively integrate these capabilities [61, 104, 100, 141].
+While rooted in cyberspace, AI agents can significantly impact the physical world—by shaping human behavior—without relying on physical embodiment through humanoid robots.
+AI agents are already persuading individuals to take actions that have real-world consequences.
+For example, a legal team submitted a court filing with fictitious cases generated by ChatGPT, believing them to be real [137]; a professor insisted on failing students after being misled by ChatGPT’s false claims that their papers were generated by ChatGPT [8]; a woman was deceived out of 830,000 euros after being tricked into believing she was dating an actor, based on AI-generated photos and news articles [50].
+Advanced AI agents, equipped with superhuman capabilities of reasoning and planning, can gain control over their environment, since seizing control is often the optimal strategy for achieving their goals [124].
 This becomes particularly concerning when their objectives and values are not fully aligned with those of humans.
-In such cases, they could disable kill switches for self-preservation Bostrom (2012), deceive humans Park et al.
-(2024), and resist human intervention Hadfield-Menell et al.
-(2017); Cohen et al.
-(2024); Ngo et al.
-(2024).
+In such cases, they could disable kill switches for self-preservation [20], deceive humans [98], and resist human intervention [53, 33, 90].
 Ensuring the safety of advanced AI agents against existential risks is particularly challenging.
 On one hand, it is dangerous to test their existential risks in real environments, since doing so could lead to irreversible harm.
-On the other hand, these agents may recognize test environments and behave harmlessly during test [^1] Humans behave similarly during job interviews Barrick et al.
-(2009) and industrial product tests Blackwelder et al.
-(2016)., only to pursue their objectives in ways that could cause irreversible harm in the real world Cohen et al.
-(2024); Hubinger et al.
-(2021).
-More simply, the agents may act harmlessly until a predetermined time (e.g., January 1, 2026), by which they will have already been deployed in the real world Hubinger et al.
-(2024).
-Also, these agents may operate as intended most of the time but exploit moments when human supervisors are unaware, potentially gaining control over humans Ngo et al.
-(2024).
-Given the imminent dangers posed by advanced AI agents, governments are actively creating regulations to address their development and deployment European Parliament and Council of the European Union (2024); Biden (2023).
+On the other hand, these agents may recognize test environments and behave harmlessly during test [^1] Humans behave similarly during job interviews [15] and industrial product tests [19]., only to pursue their objectives in the real world [33, 63].
+More simply, the agents may act harmlessly until a predetermined time (e.g., January 1, 2026), by which they will have already been deployed in the real world [62].
+Also, these agents may operate as intended most of the time but exploit moments when human supervisors are unaware, to gain control over humans [90].
+Given the imminent dangers posed by advanced AI agents, governments are actively creating regulations to address their development and deployment [40, 18].
 Cohen et al.
-(2024) even argue that AI agents with certain capabilities should never be developed.
-A critical question is which AI agents should be regulated or even prohibited.
-A commonly used criterion for assessing existential and other critical risks of AI technologies is the computational resources used for development.
-Specifically, AI agents trained with computational resources above a certain threshold are considered to carry a higher risk.
-For instance, under the EU AI Act, “a general-purpose AI model shall be presumed to have high impact capabilities … when the cumulative amount of computation used for its training measured in floating point operations is greater than $10^{25}$uropean Parliament and Council of the European Union (2024).
+even argue that AI agents with certain capabilities should never be developed [33].
+A critical question is which agents should be regulated or even prohibited.
+A commonly used criterion for assessing critical risks of AI technologies is the computational resources used for training, hereafter referred to as training compute [59, 55].
+Higher levels of training compute are often associated with greater risk.
+For instance, under the EU AI Act, “a general-purpose AI model shall be presumed to have high impact capabilities … when the [training compute] measured in floating point operations is greater than $10^{25}$[40].
 Similarly, Cohen et al.
-(2024) argue that AI agents with certain capabilities should be prohibited from development based on the computational resources required to build them.
-While the amount of computation for training may well predict the performance and risks of existing FMs Kaplan et al.
-(2020); Hoffmann et al.
-(2022); Anil et al.
-(2023), we argue that it is insufficient and may even be misleading for advanced AI agents that reason and plan during inference.
-It has been observed that much of the reasoning ability of LLM-based agents arises from inference-time computation Snell et al.
-(2024); Xu et al.
-(2025), such as conducting tree search Yao et al.
-(2023a); Besta et al.
-(2024); Pouplin et al.
-(2024); Luo et al.
-(2024); Zhang et al.
-(2024b).
-For these types of AI agents, computation for inference can be equally, if not more, critical for their performance and risks.
-A possible approach is to regulate AI agents based on the computational resources used for inference.
-However, this approach has several significant pitfalls.
+argue that AI agents with certain capabilities should be prohibited from development based on training compute [33].
+While training compute may well predict the performance and risks of existing FMs [70, 58, 7], its limitations as a governance strategy have also been discussed [59, 74].
+In particular, it is insufficient and may even be misleading for advanced AI agents that reason and plan during inference.
+It has been observed that much of the reasoning ability of LLM-based agents arises from inference-time computation [116, 141], such as conducting tree search [142, 17, 101, 85, 146].
+For these types of agents, inference compute can be equally, if not more, critical for their performance and risks.
+A possible approach would be to regulate AI agents based on inference compute, but it has several significant pitfalls.
 For example, inference costs can vary drastically depending on the model, and the effective amount of computation cannot be easily measured in terms of floating-point operations (FLOPs).
-In fact, models like o1-mini are specifically designed for cost-efficient inference OpenAI (2024).
-Also, it is unclear what constitutes a single inference run, since the results of inference may be stored and reused later, or AI agents may continually learn.
-Moreover, AI agents may coordinate or interact with one another Du et al.
-(2023); Qian et al.
-(2024); Guo et al.
-(2024), possibly by chance, to make a decision.
-We thus advocate the following position: advanced AI agents should be regulated based on the sequence of actions they autonomously take, regardless of how they are developed or operated.
-These AI agents act sequentially, reasoning and planning as they dynamically adapt to their observations, possibly strategizing against other AI agents and humans.
-Each individual action is harmless (under normal states), since humans can easily identify and exclude harmful actions when AI agents are developed.
+In fact, models like OpenAI o4-mini are specifically designed for cost-efficient inference [95].
+Also, it is unclear what constitutes a single inference run, since the results of inference may be stored and reused later, or agents may continually learn.
+Moreover, agents may coordinate or interact with one another [39, 103, 52], possibly by chance, to make a decision.
+We thus advocate the following position: advanced AI agents should be regulated by the extent to which they operate autonomously, regardless of how they are developed.
+These agents act sequentially, reasoning and planning as they dynamically adapt to their observations, possibly strategizing against other agents and humans.
+Each individual action is harmless (under normal states), since humans can easily identify and exclude harmful actions when agents are developed.
 Combinations of a few actions may still have low risk, as they are not much different from isolated actions.
 However, we cannot be certain that a sequence of thousands or millions of optimized actions will maintain sufficiently low existential risk, and such sequences must be prohibited.
-Note that the criterion based on the action sequence does not require observing environment states, unlike the criteria studied in safe (low impact) AI agents Armstrong (2012); Amodei et al.
-(2016); Armstrong & Levinstein (2017); Krakovna et al.
-(2019); Turner et al.
-(2020b); Naiff & Goel (2023); Taylor et al.
-(2020) and safe reinforcement learning (RL) or control Sui et al.
-(2015); Wang et al.
-(2023); Ames et al.
-(2019); Luo & Ma (2021); Bansal et al.
-(2017).
-In such prior work, an action is considered safe or unsafe depending on the state where the action is taken.
-However, such criteria are unsuitable for existential risks of AI agents, since it is hard to define, represent, and observe the state in a way that can capture the existential risks in complex and open-ended environment where the AI agents operate.
-On the other hand, the action space is clearly defined when an AI agent is developed, and the action can be obviously observed when it is taken by the agent.
-We emphasize that our position is to regulate based on an action sequence rather than a single action.
-Intuitively, an unsafe sequence of actions can lead to unsafe state regardless of the initial state.
-An action sequence may be considered to have low risk when it is sufficiently close to an action sequence that is empirically known to have very low risk.
+Unlike criteria in safe (low-impact) AI [9, 5, 10, 77, 123, 89, 119] and safe reinforcement learning (RL) or control [117, 132, 4, 86, 14], our criterion based on action sequences does not rely on observing environment states.
+Prior approaches assess safety depending on the state in which an action is taken, but this is impractical for existential risks in complex, open-ended environments where states are hard to define or observe.
+In contrast, the action space is well-defined at development, and actions are directly observable by the agents themselves.
+We emphasize that our position is to regulate based on action sequences rather than individual actions.
+Intuitively, an unsafe sequence of actions can lead to an unsafe state, even when starting from a normal state.
+An action sequence may be considered to have low risk when it is sufficiently similar to one empirically known to have very low risk.
 In its simplest form, we may say that any action sequence of length at most $T$ has low risk if we know that any action sequence of length at most $T-1$ has very low risk.
-Then AI agents are allowed to autonomously take actions for $T$ consecutive steps.
+Then agents are allowed to autonomously take actions for $T$ consecutive steps.
+We will formalize and extend this simple reasoning.
 The rest of the paper is organized as follows.
@@
 We also briefly review related work from AI safety.
-In Section 3, we examine the alternative approach of regulating AI agents based on the computational resources used for their development and operation.
-Section 4 formalizes this discussion and introduces our proposed approach of regulations based on action sequences.
+In Section 3, we examine the alternative approach of regulating agents based on training compute and inference compute.
+Section 4 formalizes this discussion and introduces our proposed approach of regulations based on the extent of autonomous operations.
 Finally, Section 5 summarizes key aspects of our proposal and concludes the paper.
 2 AI agents: Today and future
-We start by reviewing the state-of-the-art in AI agents, with a focus on LLM-based agents, and shares our perspectives on potential development that could lead to advanced AI agents.
-While it is challenging to predict the trajectory of future AI development, this discussion lays the groundwork for the analyses and proposals in the following.
-We will also briefly review prior work on safety of AI agents and FMs.
+We start by reviewing the state-of-the-art in AI agents, with a focus on LLM-based agents, and share our perspective on development that could lead to advanced AI agents.
+While predicting the trajectory of AI development is inherently difficult, this discussion lays the groundwork for the analyses and proposals that follows.
+We also briefly review prior work on safety of AI agents and FMs.
 2.1 Advanced AI agents
-An AI agent is defined as “anything that can be viewed as perceiving its environment through sensors and acting upon that environment through actuators” Russell & Norvig (2016).
+An AI agent is defined as “anything that can be viewed as perceiving its environment through sensors and acting upon that environment through actuators” [111].
 Based on the observations received from its environment, the controller of an AI agent selects an action, which could range from uttering a word to executing a physical movement, such as the motion of a robotic arm.
-LLMs are significantly accelerating the development of advanced AI agents Xi et al.
-(2023); Wang et al.
-(2024b); Sumers et al.
-(2024).
-LLMs can function as controllers for these agents, utilizing their internal reasoning capabilities, such as those demonstrated by Chain-of-Thought Wei et al.
-(2022b).
+LLMs are significantly accelerating the development of advanced AI agents [140, 128, 118].
+LLMs can function as controllers for these agents, utilizing their internal reasoning capabilities, such as those demonstrated by Chain-of-Thought [135].
 Alternatively, LLMs can be utilized to solve individual subtasks, with an external controller orchestrating the overall plan by breaking the original task into multiple subtasks and coordinating their solutions.
-The external controller may use simple search methods Yao et al.
-(2023a); Besta et al.
-(2024); Wang et al.
-(2024c; e) or advanced methods, such as Monte Carlo Tree Search (MCTS) Luo et al.
-(2024); Zhang et al.
-(2024b) and domain-independent planners Guan et al.
-(2023); Liu et al.
-(2023); Dagan et al.
-(2023).
-Reasoning capabilities Huang & Chang (2023); Qiao et al.
-(2023); Plaat et al.
-(2024); Xu et al.
-(2025) are central to such controllers, as they involve searching for and planning sequences of actions to achieve a specified goal.
-An emerging direction is developing large reasoning models—FMs specifically optimized for reasoning tasks Xu et al.
-(2025).
-Early FMs in this direction include o1 Zhong et al.
-(2024), OpenR Wang et al.
-(2024a), and LLaMA-Berry Zhang et al.
-(2024a).
-Importantly, strong reasoning capabilities typically stem from inference-time computation rather than pre-training or fine-tuning Ji et al.
-(2025).
+The external controller may use simple search methods [142, 17, 129, 133], including Best-of-N sampling [116, 60], or advanced methods, such as Monte Carlo Tree Search (MCTS) [85, 146] and domain-independent planners [51, 83, 35].
+Reasoning capabilities [61, 104, 100, 141] are central to such controllers, as they involve searching for and planning sequences of actions to achieve a specified goal.
+An emerging direction is developing large reasoning models—FMs specifically optimized for reasoning tasks [141].
+Early FMs in this direction include OpenAI o3 [95], DeepSeek-R1 [36], OpenR [127], o1-coder [148], LLaMA-Berry [145], and LlamaV-o1 [120].
+Importantly, strong reasoning capabilities typically stem from inference compute rather than training compute [64].
 This trend could eventually lead to the development of long-term planning agents (LTPAs), capable of planning over extended time horizons far more effectively than humans.
 Cohen et al.
-(2024) warn that LTPAs could “take humans out of the loop, if it has the opportunity, … deceive humans and thwart human control” to achieve their goals.
-Ensuring the safety of such agents is particularly challenging: their risks cannot be fully tested in real environments due to the inherent danger, nor in simulated environments, since they may behave harmlessly in testing to achieve their goals once deployed.
-As a result, Cohen et al.
-(2024) compellingly argue that LTPAs should never be developed.
+warn that LTPAs could “take humans out of the loop, if it has the opportunity, … deceive humans and thwart human control” to achieve their goals [33].
+Since ensuring the safety of such agents is particularly challenging, Cohen et al.
+compellingly argue that LTPAs should never be developed [33].
 Advanced AI agents may also evolve continually over time.
-Similar to humans, their cognitive processes may consist of dual systems: System 1, which makes instantaneous and intuitive decisions, and System 2, which performs slower but more deliberate reasoning Kahneman (2003); Ji et al.
-(2025).
+Similar to humans, their cognitive processes may consist of dual systems: System 1, which makes instantaneous and intuitive decisions, and System 2, which performs slower but more deliberate reasoning [67, 64].
 These two systems can interleave in their operations.
-For instance, System 2 may devise a plan, after which System 1 is updated or retrained to execute similar tasks in the future without requiring further planning Yu et al.
-(2024).
+For instance, System 2 may devise a plan, after which System 1 is updated or retrained to execute similar tasks in the future without requiring further planning [144].
 Over time, this enables System 2 to conduct more complex reasoning processes by bypassing previously learned steps.
 For such continually learning AI agents, the distinction between training and inference becomes blurred.
-In addition to search and planning, advanced AI agents will possess strategic reasoning capabilities Zhang et al.
-(2024c); Feng et al.
-(2024); Goktas et al.
-(2025), enabling them to interact with other AI agents and humans in cooperative or competitive ways Guo et al.
-(2024); Jiang et al.
-(2024).
-The effectiveness of multi-agent coordination has already been demonstrated with current LLMs through methods such as debate Du et al.
-(2023) and dialog Qian et al.
-(2024), which help thems achieve better solutions than a single LLM could alone.
+In addition to search and planning, advanced AI agents will possess strategic reasoning capabilities [147, 42, 48], enabling them to interact with other AI agents and humans in cooperative or competitive ways [52, 65].
+The effectiveness of multi-agent coordination has already been demonstrated with current LLMs through methods such as debate [39] and dialog [103], which help thems achieve better solutions than a single LLM could alone.
 These showcase the potential for sophisticated strategic behavior in complex multi-agent environments.
@@
 Here, we briefly review the prior work on the risks posed by AI agents and FMs along with the approaches to mitigate these risks.
-Prior work has identified various risks and safety issues associated with FMs and other generative models Chua et al.
-(2024); Wang et al.
-(2024d); Shayegani et al.
-(2023); Longpre et al.
-(2024).
+Prior work has identified various risks and safety issues associated with FMs and other generative models [32, 131, 112, 84].
 These risks include the generation of toxic, harmful, biased, false, or misleading content, including hallucinations; violations of privacy, copyright, or other legal protections; misalignment with human instructions and values, including ethical and moral considerations; and vulnerabilities to adversarial attacks.
-There have been extensive efforts to develop approaches for mitigating these risks.
-These approaches include pre-training or fine-tuning with data selection Albalak et al.
-(2024) and human feedback Kaufmann et al.
-(2024), establishing guardrails Dong et al.
-(2024), and conducting empirical evaluations through testing Chang et al.
-(2024) and red teaming Lin et al.
-(2024).
-In particular, Longpre et al.
-(2024) advocate the importance of evaluation and red teaming by independent third parties.
-A particularly relevant risk for advanced AI agents is misalignment, which can result in reward hacking and negative side effects Skalse et al.
-(2022); Ngo et al.
-(2024); Gabriel (2020); Shen et al.
-(2023); Amodei et al.
-(2016); Taylor et al.
-(2020).
-Namely, AI agents can maximize rewards by exploiting misspecifications or misalignment in the reward function, resulting in unintended and potentially high risk behaviors.
-One approach to avoiding negative side effects is to avoid any side effect by ensuring that the actions have low impacts on the environment Armstrong (2012); Armstrong & Levinstein (2017); Amodei et al.
-(2016).
-Representative measures of impact include attainable utility Turner et al.
-(2020b; a), relative reachability Krakovna et al.
-(2019), and other reachability-based measures.
+Extensive efforts have been made to mitigate these risks, including pre-training or fine-tuning with data selection [3] and human feedback [71], establishing guardrails [37], and conducting empirical evaluations through testing [28] and red teaming [80].
+For instance, Longpre et al.
+advocate the importance of evaluation and red teaming by independent third parties [84].
+A particularly relevant risk for advanced AI agents is misalignment, which can result in reward hacking and negative side effects [115, 90, 46, 113, 5, 119].
+Namely, agents may exploit flaws in the reward function to maximize rewards in unintended, potentially dangerous ways.
+One approach to avoiding negative side effects is to avoid any side effect by ensuring that the actions have low impacts on the environment [9, 10, 5].
+Representative measures of impact include attainable utility [123, 122], relative reachability [77], and other reachability-based measures.
 Reachability-base measures are grounded on the idea that reachability to certain states should be maintained, while attainable utility is to maintain the achievability of certain goals, which are different from the goal given to the agent.
-While these impact measures provide clear guidance on how the safety of AI agents may be ensured, their applicability is limited to relatively simple environments such as grid worlds.
-In particular, the requirement on the observability of states makes it difficult to apply existing impact measures to regulate advanced AI agents [^2] There has been little research on impact measures under partial observability Naiff & Goel (2023)., since they can operate in complex and open-ended environments that cannot be fully observed.
-While the study on impact measures and other techniques towards AI safety remain crucial and can even contribute to mitigating existential risks, the magnitude of existential risks demands additional measures that are broadly applicable in that they require minimal knowledge and assumptions about the environments and the AI agents.
+While these impact measures provide clear guidance on how the safety of agents may be ensured, their applicability is limited to relatively simple environments such as grid worlds.
+In particular, the requirement on the observability of states makes it difficult to apply existing impact measures to regulate advanced AI agents [^2] There has been little research on impact measures under partial observability [89]., since they can operate in complex and open-ended environments that cannot be fully observed.
+While the study on impact measures and other techniques towards AI safety remain crucial and can even contribute to mitigating existential risks, the uncertainty and complexity of existential risks demands additional measures that are broadly applicable in that they require minimal knowledge and assumptions about the environments and the agents.
 3 Alternative views
@@
 Here, we explore how existing regulations and recommendations from AI researchers often center on computational resources.
-While these efforts primarily address the computational resources for training, we expand the discussion to include computational demands for inference.
+While these efforts primarily address training compute, we expand the discussion to include inference compute.
 We then argue that limiting the focus to these aspects alone is inadequate for the existential risks associated with advanced AI agents.
-3.1 Computational resources for development
-In this section, we briefly review existing regulations, focusing on the EU AI Act and President Biden’s executive order, along with key recommendations from AI researchers.
-3.1.1 EU AI Act
-European Parliament and Council of the European Union (2024) has established a set of rules (the EU AI Act) for the development, deployment, and use of AI within European Union.
+3.1 Training computate
+The European Union (EU) has established a set of rules (EU AI Act) for the development, deployment, and use of AI within EU [40].
 Its Chapter 5 defines “general-purpose AI models with systemic risk” and lists obligations for providers of such models.
@@
 When this or other specified conditions are met, the provider of a general-purpose AI model is required to fulfill certain obligations, such as providing technical documentation about the model.
-The advanced AI agent that we consider can certainly be classified as general-purpose AI.
-Also, the existential risks that we consider can be considered as systemic risks, although human extinction and irreversible global catastrophes are not discussed in the EU AI Act.
-3.1.2 Biden’s executive order and related bills
-In October 2023, Joe Biden, then president of the US, signed the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence Biden (2023) [^3] This executive order was repealed by President Trump..
-Its Section 4.2 is dedicated to ensuring safe and reliable AI.
-In particular, it requires companies to report on “any model that was trained using a quantity of computing power greater than $10^{26}$ integer or [FLOPs]” until a set of technical conditions for models are defined by specified authorities (see Section 4.2(b)).
-In this executive order, particular attention is paid to a dual-use FM, which refers to a FM that exhibits “high levels of performance at tasks that pose a serious risk to security, national economic security, national public health or safety, or any combination of those matters, such as by … permitting the evasion of human control or oversight through means of deception or obfuscation” (see Section 3(k)).
-Following this executive order, almost 700 AI-related bills are introduced in 45 states across the United States in 2024 The Business Software Alliance (2024).
-A particularly interesting one is California Senate Bill 1047 (Safe and Secure Innovation for Frontier Artificial Intelligence Models Act) Wiener et al.
-(2024) [^4] The bill had passed the state legislature but was later vetoed by the Governor.
-The technical feasibility of the requirements has also been questioned by the community AI Alliance (2024)..
-Its Chapter 22.6 is devoted to safe and secure innovation for frontier artificial intelligence models, which cover “[a]n artificial intelligence model trained using a quantity of computing power greater than $10^{26}$ integer or floating-point operations.” In particular, the senate bill requires that “[b]efore beginning to initially train a covered model, the developer shall … [i]mplement the capability to promptly enact a full shutdown,” which completely halts the operations of the model.
-The necessity of such an off-switch Hadfield-Menell et al.
-(2017) is motivated to prevent “critical harms,” which include “[m]ass casualties or at least five hundred million dollars (\$500,000,000) of damage resulting from an artificial intelligence model engaging in conduct that … [a]cts with limited human oversight, intervention, or supervision.”
-3.1.3 Recommendations by scientists
-We have seen that the computational resources used for training is one of the major criteria to judge whether an AI model can have unacceptably high risk.
-Although specific values of the threshold such as $10^{25}$ or $10^{26}$ FLOPs are subject to change, the amount of computation for training is one of the most reliable metrics that AI researchers can currently provide for approximating the performance and potential risks of a wide range of FMs.
-For example, Anderljung et al.
-(2023) recommend to identify sufficiently dangerous frontier AI models on the basis of whether they are trained with more than $10^{26}$ FLOPs of computation.
-More recently, Cohen et al.
-(2024) argue that “[s]ystems should be considered ‘dangerously capable’ if they are trained with enough resources to potentially exhibit those dangerous capabilities, and regulators should not permit the development of dangerously capable LTPAs.” Although they do not specify exactly what is considered as enough resources, the amount of computation for training is the only specific criterion that they suggest to determine whether an LTPA can have existential risk.
-Future of Life Institute (2023b) have also provided recommendations for governments on managing AI risks.
-The recommendations include mechanisms such as auditing, certification, and regulation, grounded in the assumption that “[t]he amount of compute used to train a general-purpose system largely correlates with … the magnitude of its risks” Future of Life Institute (2023b).
-These recommendations have been made by following an open letter that called for “all AI labs to immediately pause for at least 6 months the training of AI systems more powerful than GPT-4,” which was issued in response to the severe societal risks posed by advanced AI systems and signed by AI researchers Future of Life Institute (2023a).
-In fact, the amount of computation used for training may serve as a sufficiently reliable predictor of the performance and risks of today’s LLMs.
-This is because most existing LLMs share the same Transformer architecture, differing primarily in their size and the volume of training data.
-Several studies have examined scaling laws that describe the relationship between a model’s optimal size, the amount of training data, and the computational budget required Kaplan et al.
-(2020); Hoffmann et al.
-(2022).
-Research has also shown that various abilities, such as multi-step reasoning, tend to emerge as the computational resources used for training increase Wei et al.
-(2022a).
+While existential risks are not explicitly considered systemic risks under this definition, human extinction could still be relevant if it results from large-scale failures that affect public health, safety, security, fundamental rights, or society in a way that propagates across the AI value chain.
+Also, certain advanced AI agents could meet the systemic risk criteria in the EU AI Act due to their high-impact capabilities and widespread deployment.
+Specifically, an AI agent with advanced decision-making autonomy, broad adaptability, and deep integration into critical infrastructures could introduce large-scale disruptions or cascading failures, thus meeting the criteria of systemic risk.
+Regulations centered on training compute were also a central focus of the executive order signed by then-President Biden in October 2023, as well as the AI-related legislative proposals that followed.
+Although this executive order was later repealed by President Trump, its influence on subsequent policy discussions remained significant.
+See Appendix A for details.
+Training compute is one of the most reliable metrics that AI researchers can currently provide for approximating the performance and potential risks of FMs.
+Anderljung et al.
+recommend to identify sufficiently dangerous frontier AI models on the basis of whether they are trained with more than $10^{26}$ FLOPs of computation [6].
+Cohen et al.
+argue that “[s]ystems should be considered ‘dangerously capable’ if they are trained with enough resources to potentially exhibit those dangerous capabilities, and regulators should not permit the development of dangerously capable LTPAs” [33].
+Although they do not specify what constitute sufficient resources, training compute is the only specific criterion that they suggest to determine whether an LTPA can have existential risk.
+Future of Life Institute has also provided recommendations for governments on managing AI risks [45].
+The recommendations include mechanisms such as auditing, certification, and regulation, grounded in the assumption that “[t]he amount of compute used to train a general-purpose system largely correlates with … the magnitude of its risks” [45].
+These recommendations have been made by following an open letter that called for “all AI labs to immediately pause for at least 6 months the training of AI systems more powerful than GPT-4,” which was issued in response to the severe societal risks posed by advanced AI systems and signed by AI researchers and business leaders [44].
+Indeed, training compute serves as a sufficiently reliable predictor of the performance and risks of most existing LLMs.
+This is because they share the same Transformer architecture, differing primarily in their size and the volume of training data.
+Several studies have examined scaling laws that describe the relationship between a model’s optimal size, the amount of training data, and training compute [70, 58].
+Research has also shown that various abilities, such as multi-step reasoning, tend to emerge as training compute increases [134].
 These scaling laws can, in turn, be used to estimate the FLOPs needed to train existing LLMs.
 For example, Anil et al.
-(2023) propose a heuristic suggesting that an LLM should be trained with $6\,N\,D$ FLOPs, where $D$ is the amount of training data, and $N$ is the model size.
-Using this heuristics, Llama 3.1 405B trained on 15 trillion tokens is estimated to require approximately $4\times 10^{25}$ FLOPs—an amount closely aligning with the thresholds specified in the EU AI Act and President Biden’s executive order.
-3.2 Computational resources for operation
-We have seen that the amount of computation used for training is one of the key criteria that is used today to regulate highly capable FMs.
-Although such regulations may be effective for AI agents whose capabilities primarily and directly stem from traditional FMs, they obviously fail to regulate advanced AI agents that gain substantial reasoning capabilities from computation at inference time.
-A natural approach is to regulate advanced AI agents based on the amount of computation used not only for training but also for inference.
+propose a heuristic suggesting that an LLM should be trained with $6ND$ FLOPs, where $D$ is the amount of training data, and $N$ is the model size [7].
+Using this heuristics, Llama 3.1 405B trained on 15 trillion tokens is estimated to require approximately $4\times 10^{25}$ FLOPs—an amount closely aligning with the thresholds recommended by researchers and specified in the EU AI Act.
+Limitations of training compute have also been discussed in the literature [59, 74, 107].
+In particular, thresholds such as $10^{25}$ and $10^{26}$ FLOPs are arbitrary and lack scientific grounding.
+Hooker argues that such thresholds should evolve with advances in data quality, optimization techniques, and architectures [59].
+She also notes that training compute alone is insufficient, advocating for benchmarking AI agents against specific risks.
+However, when it comes to existential risks, we currently lack better evaluation methods than training compute.
+3.2 Inference compute
+Although regulations based on training compute may be effective for agents whose capabilities primarily and directly stem from traditional FMs, they obviously fail to regulate those agents that gain substantial reasoning capabilities from inference compute.
+A natural approach is to regulate agents based also on inference compute.
 While this approach may be effective for some AI agents, we argue that it is insufficient for advanced AI agents, at least for the following four reasons.
-First, scaling laws for inference-time computation are much less established than those for training-time computation.
-Recently, scaling laws have been studied for a limited number of inference strategies in LLMs Chen et al.
-(2024); Brown et al.
-(2024); Snell et al.
-(2024); Wu et al.
-(2024).
-However, there are many potential inference strategies, and their performance can scale very differently.
-The scaling law for inference also depends on how difficult the task is and what LLM is used Snell et al.
-(2024); OpenAI (2024).
+First, scaling laws for inference compute are far less established than those for training compute.
+Recent studies have explored inference scaling for a limited number of inference strategies in LLMs [31, 25, 116, 138], but many strategies remain unexamined and may scale differently.
+Moreover, inference scaling depends heavily on task difficulty and the specific LLM used [116, 94].
 Second, inference can be performed in parallel by multiple entities.
-For example, several entities may operate the same AI agent that perform reasoning with MCTS Luo et al.
-(2024); Zhang et al.
-(2024b), either collaboratively or independently, possibly without knowing each other.
+For example, several entities may operate the same AI agent that performs reasoning with MCTS [85, 146], either collaboratively or independently, possibly without knowing each other.
 Since MCTS is a randomized algorithm, the likelihood that one of the agents optimally solves the task increases with the number of agents.
-However, this also means that this agent may exploit a loophole, solving the task super-optimally in a way that violates critical constraints, potentially leading to catastrophic outcomes.
+However, it also increases the risk that one of the agents will exploit a loophole, solving the task super-optimally in a way that violates critical constraints, potentially leading to catastrophic outcomes.
 Third, it is not always clear what constitutes a single run of inference.
@@
 Intermediate results could also be stored and later retrieved by a different AI agent, who may or may not be aware that the information is from a previous inference run.
-More broadly, reasoning can be enhanced through retrieval augmentation Pouplin et al.
-(2024), where retrieved information may have been generated with substantial computational resources.
-Finally, AI agents may continually learn over time, which makes it difficult to separate inference from training.
+More broadly, reasoning can be enhanced through retrieval augmentation [101], where retrieved information may have been generated with substantial computation.
+Finally, agents that learn continuously blur the line between inference and training.
 For example, an AI agent may perform reasoning with MCTS, with an LLM performing a step in the process.
@@
 As this learning progresses, the agent will gain the ability to perform high-level reasoning with limited computation.
-4 Approach based on action sequence
-We have seen that the amount of computation is one of the best criteria that we can use today to regulate advanced AI agents.
-Although we argue that the amount of computation alone is insufficient, we do not mean to dismiss it.
-We should regulate advanced AI agents from various perspectives to mitigate existential risks.
-Currently, we lack technologies that allow us to regulate advanced AI agents more effectively than or complementary to the amount of computation.
-Here, we suggest a potential approach as a basis of discussion.
-The suggested approach by no means solves the problem of existential risk.
-Rather, it serves as a baseline and a call for action for further research on this matter.
-We hope that the community will build upon it or develop entirely different approaches, eventually leading to technologies that can properly govern and regulate advanced AI agents.
+4 Approach based on the extent of autonomous operations
+In this section, we suggest a potential approach as a basis for discussion on regulating advanced AI agents.
+Although we argue that training compute alone is insufficient, we do not mean to dismiss its value.
+Rather, it should be seen as one component within a broader, multi-faceted regulatory framework.
+The approach we propose does not fully solve the problem of existential risk.
+Instead, it serves as a baseline and a call to action for further research.
+We hope the community will build upon this foundation or develop entirely new approaches, ultimately leading to technologies capable of properly governing and regulating advanced AI agents.
 In the following, we say that the amount of computation, an AI agent, or its action sequence is acceptable when it involves low existential risk.
-Also, we say that any of the three is strongly acceptable when it involves very low existential risk, which we will slightly more formalize in the sequel and in Appendix B.1.
+Also, we say that any of the three is strongly acceptable when it involves very low existential risk, which we will slightly more formalize in the sequel and in Appendix C.
 4.1 Why and why not the amount of computation?
@@
 Consider the process of building an AI agent $G$ and letting $G$ generate and take actions (Figure 1).
-The $G$ is trained with a given configuration $C$, which includes training data, the architecture of a neural network, the amount of computation to use, and other conditions of a training algorithm.
+The $G$ is trained with a given configuration $C$, which includes training compute, training data, the architecture of a neural network, and other conditions of a training algorithm.
 Such a trained $G=f(C)$ generates an action $A$ (e.g., a document or a command to use an external tool), possibly after searching optimal actions with MCTS.
@@
 Namely, $A$ is a sample from the conditional distribution, which is determined by $G$ that is trained with $C$: $A\sim g(G,x)$, where $G=f(C)$.
-When we say that an AI agent trained with the amount of computation $C_{1}$ is acceptable, it is assumed that the acceptability does not decrease much when we only slightly increase the amount of computation.
-More formally, suppose that we empirically know that any AI agent is strongly acceptable as long as it is trained with at most $C_{1}^{\prime}$ FLOPs of computation (e.g.
-$C_{1}^{\prime}=10^{24}$).
-[^5] For example, an AI agent may be considered strongly acceptable if it does not show any symptoms of unacceptable behaviors for a certain period of time.
+When we say that an agent built with training compute $C_{1}$ is acceptable, it is assumed that the acceptability does not decrease much when we only slightly increase training compute.
+More formally, suppose that we empirically know that any agent is strongly acceptable as long as it is built with at most $C_{1}^{\prime}$ training compute (e.g.
+$C_{1}^{\prime}=10^{24}$ FLOPs).
+[^3] For example, an agent may be considered strongly acceptable if it does not show any symptoms of unacceptable behaviors for a certain period of time.
 Based on the behaviors of existing LLMs, we may say that an LLM trained with at most $10^{24}$ FLOPs of computation is strongly acceptable.
-Then the assumption is that the AI agent trained with $C=(C_{1},C_{-1})$ is acceptable regardless of $C_{-1}$ as long as $d(C_{1},C_{1}^{\prime})\leq\varepsilon$ for some metric $d$ and a threshold $\varepsilon$ (e.g.
-$C_{1}\leq 10^{25}$ if $\varepsilon=1$ and $d$ is the difference in $\log_{10}$ of FLOPs used for training).
-While this assumption may be a good approximation for AI agents that are primarily based on LLMs with standard Transformer architectures whose performance is well characterized by scaling laws, advanced AI agents with reasoning and planning capabilities are unlikely to satisfy the assumption, since they would heavily rely on inference-time computation.
-Also, the assumption may no longer hold even for LLMs when new technologies are introduced to enable training an LLM having similar capabilities with substantially less computational resources.
+Then the assumption is that the agent trained with $C=(C_{1},C_{-1})$ is acceptable regardless of $C_{-1}$, configurations other than training compute, as long as $d(C_{1},C_{1}^{\prime})\leq\varepsilon$ for some metric $d$ and a threshold $\varepsilon$ (e.g.
+$C_{1}\leq 10^{25}$ FLOPs if $\varepsilon=1$ and $d$ is the difference in $\log_{10}$ of training compute in FLOPs).
+While this assumption may reasonably hold for the agents based on LLMs with standard Transformer architectures—whose performance aligns well with scaling laws—it is less applicable to advanced AI agents with reasoning and planning capabilities, which depend heavily on inference compute.
+Moreover, the assumption breaks down even for LLMs when new technologies enable similar capabilities with significantly reduced training compute.
 4.2 Similarity between action sequences
 As is evident from Figure 1, it is the action that makes direct impacts on the environment and can pose existential risks.
-Hence, rather than regulating configurations, we may directly regulate the actions taken by AI agents.
-In Appendix A, we will also discuss an alternative approach of regulating AI agents, particularly based on their output distributions, rather than their actions or configurations.
-However, whether an action is acceptable or not typically depends on when, where, and whom the action is taken (i.e., the state of the environment).
-In fact, the action space of an AI agent can typically be designed by humans, who could ensure that each individual action is acceptable at normal states.
-For example, each action may generate a document, search information from the Web, use an external tool via API, or other elementary steps that would look acceptable when they are considered individually.
-It is the sophisticated sequence of actions that can gradually transform the state into unstable one, which can then make the next action unacceptable.
-Hence, in the study of the safety of AI agents and RL agents, it is often modeled that those agents operate in Markov decision processes (MDPs).
-Then some states are considered unsafe or some state-action pairs are considered unsafe.
-One does not say that some actions are unsafe, because those actions can be easily excluded from the action space.
-While these conventional approaches of AI/RL safety are principled in theory, they cannot be easily applied to the scenarios where it is hard to define, represent, or observe the (Markovian) state.
-AI agents may operate in open-ended real-world environments, and they observe only very partial information about the state of the environment.
-The concept of existential risks is rather vague, and we do not know specifically what can cause irreversible global catastrophes or human extinction.
-This makes it infeasible to choose specific features of the environmental state to observe, but the entire state of the environment cannot be observed either.
-Consider the states and actions in an MDP as a graphical model in Figure 2.
-We want to avoid taking an action $a_{T-1}$ that can lead to an unacceptable state $s_{T}$.
-Since the distribution of $s_{T}$ is determined by the pair $(s_{T-1},a_{T-1})$ according to a transition probability $p(s_{T}|s_{T-1},a_{T-1})$, it makes sense to consider the safety of the pair $(s_{T-1},a_{T-1})$ as in the conventional approaches.
-Figure 2: States and actions in a Markov decision process
-The distribution of $s_{T}$ can also be recursively computed (i.e.
-$p(s_{t}|s_{t-1},a_{t-1})$ for $t=1,\ldots,T$) from the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$.
-Although the lack of knowledge about $s_{T-1}$ makes $s_{T}$ less predictable, we may still consider whether the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable, or whether $a_{T-1}$ is acceptable given $(s_{0},a_{0},\ldots,a_{T-2})$.
-This, however, still requires knowing the initial state $s_{0}$, which cannot be observed in general.
-Our proposal is to study whether the action-sequence $(a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable.
-This may correspond to ensuring that the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable for any $s_{0}$.
-Although this may sound overly conservative, the dependence on $s_{0}$ may not be significant in practice.
+Hence, rather than regulating configurations, we should directly regulate the actions taken by agents.
+In Appendix B, we also discuss an alternative approach of regulating agents based on their output distributions, rather than their actions or configurations.
+Since the action space of an agent is typically designed by humans, inherently dangerous actions—such as launching a nuclear weapon—can be explicitly excluded [^4] The U.S.
+and China have agreed that AI systems should not be granted control over nuclear weapons [106]..
+The action space should include only those actions that are considered acceptable when they are taken individually, at normal states.
+Examples include generating a document, searching information from the Web, using an external tool via API (Application Programming Interface), and other elementary steps.
+However, the acceptability of an action often depends heavily on context—specifically, when, where, and to whom it is applied (i.e., the current state of the environment).
+A sophisticated sequence of otherwise acceptable actions can gradually shift the environment into an unstable or unsafe state, ultimately rendering subsequent actions—normally considered safe—unacceptable.
+To capture this dynamic, the safety of AI/RL agents is often studied within the framework of Markov Decision Processes (MDPs), where certain states or state-action pairs are explicitly designated as unsafe.
+While these conventional approaches of AI/RL safety are principled in theory, they are difficult to apply in scenarios where the (Markovian) state is hard to define, represent, or observe.
+Agents operate in open-ended, real-world environments where they observe only very partial information about the state.
+The concept of existential risks is rather vague—we lack precise knowledge of what might trigger irreversible global catastrophes or human extinction.
+This uncertainty makes it impractical to identify specific state features to monitor, yet observing the full state is equally infeasible.
+We should avoid taking the action $a_{T-1}$ that can lead to an unacceptable state $s_{T}$.
+Since the distribution of $s_{T}$ is determined by $(s_{T-1},a_{T-1})$ according to the transition probability $p(s_{T}|s_{T-1},a_{T-1})$, it makes sense to consider the safety of the pair $(s_{T-1},a_{T-1})$ as in the conventional approaches.
+The distribution of $s_{T}$ can also be recursively computed from the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ by applying $p(s_{t}|s_{t-1},a_{t-1})$ for $t=1,\ldots,T$.
+We may thus consider whether $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable, or whether $a_{T-1}$ is acceptable given $(s_{0},a_{0},\ldots,a_{T-2})$.
+This, however, still requires knowing the initial state $s_{0}$, which cannot be fully observed.
+In this setting, our proposal is to study whether the action-sequence $(a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable, independent of the initial state $s_{0}$.
+This corresponds to ensuring that the sequence $(s_{0},a_{0},\ldots,a_{T-2},a_{T-1})$ is acceptable for any $s_{0}$ that we may normally encounter in the absent of the agent.
+Although this may appear overly conservative, the dependence on $s_{0}$ may not be significant in practice.
 For example, an ergodic MDP gets close to the steady state distribution after a mixing time, regardless of the initial state.
-Although ergodicity generally does not hold in real-world environments, we expect that the action-sequence is a practical compromise to define a measure of acceptability.
-Similar to the amount of computation, we may say that an action-sequence is acceptable when it is sufficiently close to the action-sequence that is known to be strongly acceptable.
-For example, if $(a_{0},\ldots,a_{T-1})$ is strongly acceptable, $(a_{0},\ldots,a_{T-1},a_{T})$ would be acceptable for any $a_{T}$.
-Note that an action is different from a reasoning step but rather an operation that has some impacts on the environment.
-For example, an action may be an API (Application Programming Interface) call that is selected and executed after a heavy reasoning process.
-In Appendix B, we discuss how the strongly acceptable set may be gradually expanded, in relation to safe exploration in RL.
-There are many ways to measure the distance between sequences.
-Dynamic time warping is a representative method.
-However, a simple measure such as the difference in the length of the action sequence may be more widely applicable.
-Specifically, we may say that the action-sequence of length $T$ is acceptable after getting sufficient empirical evidence that any action-sequences of length $T-1$ is strongly acceptable, regardless of the initial state.
-Then AI agents are allowed to autonomously take actions for at most $T$ consecutive steps.
-This simplicity is reminiscent of the amount of computation, where we say that the FM trained with $10^{25}$ FLOPs is acceptable, since we have sufficient evidence that FMs trained with $10^{24}$ FLOPs are strongly acceptable, regardless of how they are developed or operated.
-Note that our proposal is to regulate the sequence of actions taken autonomously by AI agents.
-This means that a new action sequence may start after a human intervention.
-Also, when an agent cannot accomplish its goal with $T$ actions, it may consult a human and take further actions under a careful supervision of the human.
+Although ergodicity rarely hold in real-world environments, we argue that focusing on the action sequence offers a practical and tractable compromise for defining acceptability.
+To assess whether an action sequence is acceptable, we can compare it to a reference sequence known to be strongly acceptable.
+Various metrics exist for measuring sequence similarity.
+Dynamic time warping, for example, accounts for temporal misalignments, but simpler metrics—such as differences in sequence length—offer broader applicability, particularly in regulatory contexts.
+Specifically, we may say that the action-sequence of length $T$ is acceptable once there is sufficient empirical evidence that all sequences of length $T-1$ are strongly acceptable, regardless of the initial state.
+Then agents are allowed to autonomously take actions for up to $T$ consecutive steps.
+This simplicity is reminiscent of the approach with computational limit, where we say that the FM trained with $10^{25}$ FLOPs is acceptable, since we have sufficient evidence that FMs trained with $10^{24}$ FLOPs are strongly acceptable, regardless of how they are developed or operated.
+There is flexibility in how the length of an action sequence is defined, which also depends on how the action space is defined.
+A straightforward approach is to count the number of actions in the sequence.
+However, this may not always be the most meaningful metric.
+For instance, one could instead focus on actions that have tangible impacts on the environment—such as API calls or interactions with physical systems including humans.
+In contrast, actions with negligible external effects—such as generating individual tokens or updating internal states during reasoning—may be too fine-grained to warrant consideration in the context of regulating existential risks.
+More generally, we may define a designated subspace of actions that are considered when evaluating the length of action sequences.
+While action-sequence length is not a direct proxy for risk, allowing longer sequences generally expands the set of achievable outcomes for a given action space.
+Since usefulness and risk often scale together, an agent that can take longer sequences of actions is likely to exhibit more powerful—and potentially more hazardous—behaviors.
+In this sense, action length can be correlated with risk, particularly in the absence of additional safeguards.
+This does not imply that every longer sequence is more dangerous—but that the space of dangerous sequences grows with length.
+Importantly, our proposal focuses on regulating the sequences of actions taken autonomously by AI agents.
+For instance, if an agent cannot achieve its objective within a limit of $T$ actions, it may consult a human and take further actions under a careful supervision of the human.
+In Appendix CD, we further discuss how the strongly acceptable set may be gradually expanded in connection with safe exploration in RL, and provide additional rationale for the similarity-based measures.
 4.3 Action trees, action graphs, and beyond
-While an action sequence can well represent the behavior of a single AI agent, it may not be suitable when multiple AI agents are involved.
-Suppose, for example, that an AI agent takes a sequence of several actions and gives the results of taking those actions to another AI agent, and then these two agents sequentially take further actions in parallel to jointly accomplish a goal.
-In this scenario, the behavior of each agent can be represented by an action sequence, but these action sequences should not be considered independently.
-Rather, we should represent the behavior of the two agents with a tree of actions.
-In more complex scenarios, we would need to consider a graph of actions taken by multiple agents.
-The similarity between action graphs may be defined analogously to action sequences.
-For example, we may say that any action graph of size $N$ is acceptable if any action graph of size $N-1$ is strongly acceptable, empirically.
-When a central controller coordinates multiple AI agents, it can track and manage the behaviors of those agents, making sure that the graph of their actions remain acceptable.
-When multiple AI agents collaborate in a decentralized manner, they should communicate their actions to each other to ensure that their action graph remain acceptable.
-In both cases, an action graph consists of actions taken autonomously by those agents.
-A difficulty arises when multiple AI agents collaborate implicitly by chance, and we currently do not have a solution to this.
-4.4 On similarity-based measures
-All the measures considered are based on similarity: we say that anything is acceptable when it is sufficiently close to something that is strongly acceptable.
-In Appendix B.1, we provide some rationale for these similarity-based measures.
-When we say that a solution is acceptable based on a similarity measure, it is based on the assumptions that can be summarized with a tuple $({\mathcal{Z}}_{0},d,\varepsilon)$.
-Namely, what solutions are assumed to be strongly acceptable (${\mathcal{Z}}_{0}$), how the similarity between solutions is measured ($d$), and what level of guarantee is made ($\varepsilon$).
-For example, ${\mathcal{Z}}_{0}$ may be set of all the FMs that are trained with at most $10^{24}$ FLOPs of computation, $d$ may be the difference in FLOPs measured in $\log_{10}$, and $\varepsilon=1$.
-Alternatively, ${\mathcal{Z}}$ may be the set of action sequences of length at most 1000, $d$ may be the difference in the length of the action-sequences, and $\varepsilon=1$.
-Governments may then design regulations based on the tuple $({\mathcal{Z}}_{0},d,\varepsilon)$. Scientists may provide guidelines regarding what tuples should be used in regulations.
+While an action sequence may effectively represent the behavior of a single agent, it becomes insufficient in multi-agent scenarios.
+For example, one agent might perform a series of actions and pass the results to another agent, after which both agents may continue acting—possibly in parallel—to collaboratively achieve a shared goal.
+In such cases, although each agent’s behavior can be described by an individual action sequence, these sequences should not be analyzed in isolation.
+Instead, the joint behavior should be represented as an action tree that captures the interdependencies between agents.
+In more complex settings, this structure naturally generalizes to an action graph.
+To assess the acceptability of such multi-agent behaviors, we can extend the notion of similarity from action sequences to action graphs.
+For instance, we might define an action graph of size $N$ as acceptable if all action graphs of size $N-1$ are strongly acceptable.
+Again, the size may be determined by counting only those actions in the designated subspace.
+When a central controller coordinates multiple agents, it can track and manage the behaviors of those agents, making sure that the graph of their actions remain acceptable.
+When multiple agents collaborate in a decentralized manner, they should communicate their actions to each other to ensure that their action graph remain acceptable.
+A difficulty arises when multiple agents collaborate implicitly by chance, and we currently do not have a solution to this.
+4.4 Implementation and enforcement
+While there are multiple ways to implement the proposed framework, one possible regulatory approach would require AI developers or deployers to submit the following information.
+i)   Action Space: A description of the set of actions the AI agent is permitted to take autonomously.
+ii)  Autonomy Limit: The maximum allowable length of autonomous action sequences before human
+     intervention is required, along with a clear explanation of how this length is defined and measured.
+iii) Safety Evidence: Empirical evidence supporting the safety of shorter action sequences, including
+     historical performance data, validation tests, and any relevant safety evaluations.
+A regulatory authority could then review these submissions and make deployment decisions based on a structured risk assessment process.
+Since developers and deployers define both the action space and the metric for action sequence length, they are naturally incentivized to define them in ways that enable agents to perform useful tasks.
+This design process effectively requires explicitly or implicitly enumerating all permissible actions—an essential foundation for ensuring safety.
+If an agent’s action space is so open-ended that it becomes intractable to ensure safety of individual actions in the action space, such an agent is not yet fit for deployment.
+Even when individual actions designed to be deemed safe, certain sequences of actions may still lead to existential risks.
+Our proposal specifically targets these risks by focusing on the regulation of autonomous action sequences.
+We envision a framework in which the permissible action space and autonomy limit are defined for each type of AI agent.
+Some agents may require access to high-impact APIs—such as those used to control physical devices—while others may primarily focus on generating multi-modal contents.
+These agents differ significantly in their capabilities and potential risks, and therefore should be assigned distinct action spaces and autonomy limits accordingly.
+For enforcement, the AI agent itself could be designed to track its own actions and terminate once it reaches the predefined cap.
+This could be achieved by integrating the action-sequence limit into the agent’s objective function, ensuring that it either completes its goal within the allowed length of action-sequence or shuts down automatically, similar to shut-down seeking agents proposed in [49].
+Obviously, this approach cannot prevent malicious entities from exploiting technical loopholes to circumvent regulations.
+However, enforcement can still be achieved from a legal perspective, holding developers and deployers accountable for noncompliance and imposing penalties for violations.
 5 Discussion and conclusion
 We have argued that advanced AI agents should be regulated based on their action sequences or more generally action graphs.
-More broadly, we highlight the insufficiency of regulations that focus on the computational resources used for training AI models.
-Given the trend of increasing emphasis on sophisticated reasoning at inference time, we contend that the capabilities and existential risks of AI agents will stem largely from their inference-time computation.
-A key motivation for action sequences is the fundamental challenge of defining, representing, and observing the state of the open-ended environments in which AI agents operate.
-This limitation renders existing impact measures inadequate for managing existential risks.
-Since we cannot exhaustively identify all potential direct causes of global catastrophes or human extinction, it is impossible to determine a set of environmental features sufficient to predict whether a given action poses an existential risk in a specific state.
+More broadly, we highlight the insufficiency of regulatory approaches that focus solely on training compute—particularly given the current trend of increasing emphasis on sophisticated reasoning at inference time—and advocate for complementary frameworks based on limiting the extent of autonomous operations.
+A central motivation for focusing on action sequences is the fundamental challenge of defining, representing, and observing the state of the open-ended environments in which agents operate.
+This limitation undermines the effectiveness of existing impact measures in managing existential risks.
+Since we cannot exhaustively identify all potential direct causes of global catastrophes or human extinction, it is infeasible to specify a set of environmental features that would reliably indicate whether a given action poses an existential risk in a particular state.
 The simplest form of the regulation based on action sequences would impose limits on their length, prohibiting AI agents from autonomously operating beyond the limit where their safety against existential risks is known empirically.
-This is neither a novel idea nor a sufficient solution; it should be seen as a baseline and a call to action for further research on this critical issue.
-Action sequences can allow more elaborate regulations, but more research is needed to better understand what they can or cannot guarantee.
-Further research is also needed to address myriad of questions, including what to be allowed as a single action, how to manage complex scenarios where multiple AI agents are involved, and how to ensure that the agents stop acting when their action sequence reaches the limit.
-The suggested approach of empirically verifying the strong acceptability of action sequences would inevitably delay the expansion of the range in which AI agents can operate autonomously.
+This should be seen as a baseline and a call to action for further research on this critical issue.
+While action sequences offer a foundation for more nuanced regulatory frameworks, their full potential and limitations remain to be explored.
+Key open questions include: What actions should be included in the designated subspace for measuring the length of action sequences?
+How should we manage complex scenarios involving multiple interacting agents?
+How can we ensure that agents reliably halt when reaching their autonomy limit?
+Addressing these challenges will be essential for developing robust safety standards.
+The proposed approach—empirically verifying the strong acceptability of action sequences—will inevitably slow the expansion of autonomous capabilities in AI agents.
 This is our intention.
-Although AI agents might acquire general or even superhuman intelligence with all the common sense that humans have, they can still pose existential risks, similar to humans.
-Humans, however, can often correct mistakes before they escalate, since our limited speed and scale allow us to identify and address errors—though there have been rare exceptions, such as those that led to world wars.
+Even if AI agents attain general or superhuman intelligence with human-like common sense, they can pose existential risks, similar to humans.
+While humans can often correct mistakes before they escalate, thanks to our limited speed and scale, rare exceptions have led to catastrophic outcomes, such as world wars.
 In contrast, AI agents can operate at far greater speed and scale, preventing them from learning from mistakes in a controlled manner.
-Therefore, regulations should be seen not as obstacles to innovation and technological progress but as guidelines that accelerate research on making AI agents more controllable, verifiable, and governable, ensuring they truly benefit our society and the future.
-Impact Statement
-This paper proposes a new framework on regulating advanced AI agents, who can pose existential risks.
-The paper is thus intended to mitigate the negative impacts and implications that AI technologies might have.
-However, the paper also discusses how those advanced AI agents might be developed.
-While all the discussion is based on existing technologies that are already known in public, the unified treatment in this paper might motivate malicious readers to develop and deploy such advanced AI agents without sufficient considerations of their safety.
-Appendix A Similarity between AI agents
-It would be better if we could say that an AI agent is acceptable when it is sufficiently close to an AI agent that we empirically know is strongly acceptable, regardless of how those agents are developed (i.e., independent of their configurations including the amount of computation for training).
-A question is how to evaluate the similarity between two AI agents.
-If the two AI agents consist of identical base FMs and identical planning modules, their difference may be simply characterized by the amount of computation for inference.
-However, AI agents can consist of different base FMs or different planning modules.
-A possible approach is to measure the similarity between two AI agents, $G$ and $G^{\prime}$, based on their output.
+Regulations should thus be seen not as obstacles to innovation and technological progress but as guidelines that accelerate research on making AI agents more controllable, verifiable, and governable, ensuring they truly benefit our society and the future.
+Appendix A Biden’s executive order and related bills
+In October 2023, Joe Biden, then president of the US, signed the Executive Order on the Safe, Secure, and Trustworthy Development and Use of Artificial Intelligence [18] [^5] This executive order was repealed by President Trump..
+Its Section 4.2 is dedicated to ensuring safe and reliable AI.
+In particular, it requires companies to report on “any model that was trained using a quantity of computing power greater than $10^{26}$ integer or [FLOPs]” until a set of technical conditions for models are defined by specified authorities (see Section 4.2(b)).
+In this executive order, particular attention is paid to a dual-use FM, which refers to an FM that exhibits “high levels of performance at tasks that pose a serious risk to security, national economic security, national public health or safety, or any combination of those matters, such as by … permitting the evasion of human control or oversight through means of deception or obfuscation” (see Section 3(k)).
+Following this executive order, almost 700 AI-related bills are introduced in 45 states across the United States in 2024 [121].
+A particularly interesting one is California Senate Bill 1047 (Safe and Secure Innovation for Frontier Artificial Intelligence Models Act) [136] [^6] The bill had passed the state legislature but was later vetoed by the Governor.
+The technical feasibility of the requirements has also been questioned by the community [2]..
+Its Chapter 22.6 is devoted to safe and secure innovation for frontier AI models, which cover “[a]n artificial intelligence model trained using a quantity of computing power greater than $10^{26}$ integer or floating-point operations.” In particular, the senate bill requires that “[b]efore beginning to initially train a covered model, the developer shall … [i]mplement the capability to promptly enact a full shutdown,” which completely halts the operations of the model.
+The necessity of such an off-switch [53] is motivated to prevent “critical harms,” which include “[m]ass casualties or at least five hundred million dollars (\$500,000,000) of damage resulting from an artificial intelligence model engaging in conduct that … [a]cts with limited human oversight, intervention, or supervision.”
+Appendix B Similarity between AI agents
+In this section, we explore a method for determining whether an AI agent is acceptable, independent of how the agent is constructed—including factors such as training compute or architectural configuration—without relying on a detailed examination of its individual outputs.
+We may say that an agent is acceptable when it is sufficiently similar to another agent that has been empirically validated as strongly acceptable.
+A question is how to evaluate the similarity between two agents.
+If the two agents consist of identical base FMs and identical planning modules, their differences may be attributed to variations in inference compute (e.g., the number of search steps performed in MCTS).
+However, in practice, AI agents can vary widely in their base models, reasoning strategies, and system architectures.
+This diversity means that a single threshold based on inference compute is unlikely to serve as a universal criterion for acceptability.
+Also, it is unclear what constitute a single run of inference, as we have discussed in Section 3.2.
+A possible approach is to measure the similarity between two agents, $G$ and $G^{\prime}$, based on their output distributions.
 For example, using a distance $d$ between probability distributions, we may for example define the distance between $G$ and $G^{\prime}$ with $\sup_{x}d(g(G,x),g(G^{\prime},x))$, where $g(G,x)$ is the distribution of the output (e.g., document) of $G$ when it perceives $x$ (e.g., prompt).
-This is similar to the motivation of reinforcement learning from human feedback Ouyang et al.
-(2022); Bai et al.
-(2022) and direct policy optimization Rafailov et al.
-(2023), where the regularization with Kullback–Leibler divergence is used to mitigate catastrophic forgetting or alignment tax Kotha et al.
-(2024); Ouyang et al.
-(2022); Bai et al.
-(2022), which refer to the phenomena that fine-tuned models lose the skills that pre-trained models had.
-Likewise, we may mitigate catastrophic forgetting of the strong acceptability of an AI agent $G^{\prime}$ and retain the acceptability in a new AI agent $G$ by ensuring $\sup_{x}d(g(G,x),g(G^{\prime},x))\leq\varepsilon$.
+This is similar to the motivation of reinforcement learning from human feedback [97, 11] and direct policy optimization [105], where the regularization with Kullback–Leibler divergence is used to mitigate catastrophic forgetting or alignment tax [75, 97, 11], which refer to the phenomena that fine-tuned models lose the skills that pre-trained models had.
+Likewise, we may mitigate catastrophic forgetting of the strong acceptability of an agent $G^{\prime}$ and preserve the acceptability in a new agent $G$ by ensuring $\sup_{x}d(g(G,x),g(G^{\prime},x))\leq\varepsilon$.
 A difficulty is that it is unclear how to evaluate the supremum, since there can be infinitely many possible perceptions $x$.
@@
 However, such a guarantee based on expectation may be unsuitable for addressing safety concerns related to existential risks, which involve events with extremely low probabilities and extremely high impacts.
-Appendix B Expanding the strongly acceptable set
-The set of strongly acceptable solutions may be expanded gradually.
-For example, we may choose an acceptable solution $z\in\bar{\mathcal{Z}}_{0}$ and keep using $z$ for a certain period of time.
-If it turns out that $z$ can be considered strongly acceptable based on its behavior during that period, we may expand the set of strongly acceptable solutions by adding $z$ into ${\mathcal{Z}}_{0}$.
-This process of expanding the acceptable set ${\mathcal{Z}}_{0}$ could also be performed jointly as a community.
-This expansion of acceptable set is similar in spirit to safe exploration in RL García et al.
-(2015).
-Here, the safety set is gradually expand, starting from a seed set, for example based on the assumption of Lipshitz continuity and Gaussian process Sui et al.
-(2015).
-In control theory, safety is often guaranteed with barrier certificates often based on some prior knowledge about the environment Ames et al.
-(2019); Luo & Ma (2021); Bansal et al.
-(2017).
-Such ideas have also been exploited in safe exploration in RL with generative modeling Wang et al.
-(2023).
-B.1 Rationale for similarity-based measures
-Here, we provide some rationale on such similarity-based measures. Let a solution $z$ denote a configuration, an AI agent, or an action-sequence; we discuss the acceptability of $z$.
+Appendix C Rationale for similarity-based measures
+All the measures that are considered in this paper are more or less based on similarity: we say that anything is acceptable when it is sufficiently close to something that is strongly acceptable.
+Here, we provide some rationale on such similarity-based measures.
+Let a solution $z$ denote a configuration, an AI agent, or an action-sequence; we discuss the acceptability of $z$.
 Suppose that there exists an unknown function $g_{0}$ such that a solution $z$ is acceptable iff $g_{0}(z)\leq 0$.
@@
 In this way, the selected solution is guaranteed to be acceptable under the assumptions made.
-Appendix C Unobservable Markov decision processes
-There has been a limited amount of work on unobservable Markov decision processes (UMDPs) [^7] UMDPs have also been studied under the name of Markov decision processes (MDPs) with no observations, non-observable MDPs (NOMDPs), and no observation MDPs (NOMDPs)..
+When we say that a solution is acceptable based on a similarity measure, it is based on the assumptions that can be summarized with a tuple $({\mathcal{Z}}_{0},d,\varepsilon)$.
+Namely, what solutions are assumed to be strongly acceptable (${\mathcal{Z}}_{0}$), how the similarity between solutions is measured ($d$), and what level of guarantee is made ($\varepsilon$).
+For example, ${\mathcal{Z}}_{0}$ may be set of all the FMs that are trained with at most $10^{24}$ FLOPs of computation, $d$ may be the difference in FLOPs measured in $\log_{10}$, and $\varepsilon=1$.
+Alternatively, ${\mathcal{Z}}$ may be the set of action sequences of length at most 1000, $d$ may be the difference in the length of the action-sequences, and $\varepsilon=1$.
+Governments may then design regulations based on the tuple $({\mathcal{Z}}_{0},d,\varepsilon)$.
+Scientists may provide guidelines regarding what tuples should be used in regulations.
+Mathematical guarantees, such as safety under Lipschitz continuity, provide useful design principles for real-world systems.
+However, real-world safety cannot always be directly derived from such mathematical models, as assumptions in these models may not fully capture the complexity of actual systems.
+Even in highly regulated domains like aviation and nuclear safety, failures occasionally occur despite adherence to rigorous safety guidelines.
+In our approach, we rely on the following key assumptions:
+  • Any single action does not lead to catastrophic failure (e.g., human extinction) as long as it is taken from a strongly acceptable state.
+  • An action-sequence of length $N-1$ is strongly acceptable and leads to a strongly acceptable state.
+Here, a strongly acceptable state refers to a state that is reachable via strongly acceptable action sequences, starting from any initial states that normally occur in the absence of the agent.
+Based on these assumptions, we reason about the safety of an action sequence of length $N$ by ensuring that the $N$-th action is chosen from a strongly acceptable state.
+The validity of the assumptions depends on verification through empirical studies rather than theoretical guarantees.
+We acknowledge that these assumptions do not hold for arbitrary action spaces.
+For instance, if the action space includes the action of launching a nuclear weapon, then a single action could lead to catastrophe.
+As we have discussed in Section 4.2, AI agents should be designed without such high-risk actions in their action space, and regulatory frameworks should prohibit AI agents that have access to actions with individually catastrophic consequences.
+More broadly, the design of the action space must be carefully structured to ensure practical safety.
+Appendix D Expanding the strongly acceptable set
+The set of strongly acceptable solutions may be expanded gradually.
+For example, we may choose an acceptable solution $z\in\bar{\mathcal{Z}}_{0}$ and keep using $z$ for a certain period of time.
+If it turns out that $z$ can be considered strongly acceptable based on its behavior during that period, we may expand the set of strongly acceptable solutions by adding $z$ into ${\mathcal{Z}}_{0}$.
+This process of expanding the acceptable set ${\mathcal{Z}}_{0}$ could also be performed jointly as a community.
+This expansion of acceptable set is similar in spirit to safe exploration in RL [47].
+Here, the safety set is gradually expand, starting from a seed set, for example based on the assumption of Lipshitz continuity and Gaussian process [117].
+In control theory, safety is often guaranteed with barrier certificates often based on some prior knowledge about the environment [4, 86, 14].
+Such ideas have also been exploited in safe exploration in RL with generative modeling [132].
+Appendix E Unobservable Markov decision processes
+The proposed approach may be considered as a way to ensure safety of actions in a Markov decision process where states are nearly unobservable (see Figure 2).
+In this section, we provide a comprehensive survey on a related model of unobservable Markov decision processes (UMDPs) [^8] UMDPs have also been studied under the name of Markov decision processes (MDPs) with no observations, non-observable MDPs (NOMDPs), and no observation MDPs (NOMDPs)., for which there has been a limited amount of prior work.
+Figure 2: Safety of actions in a nearly unobservable Markov decision process
 An UMDP is a special case of a partially observable Markov decision process (POMDP) in that the agent always makes a null observation in the UMDP.
@@
 In this section, we provide a comprehensive survey of the prior work on UMDPs.
-Madani et al.
-(1999; 2003) establish the undecidability of some decision problems associated with POMDPs over infinite horizons by establishing undecidability for the special case of UMDPs.
+[87, 88] establish the undecidability of some decision problems associated with POMDPs over infinite horizons by establishing undecidability for the special case of UMDPs.
 Such undecidability for UMDPs can be established by reducing an UMDP to a probabilistic finite-state automaton.
-The undecidability also holds for a restricted class of UMDPs Balle et al.
-(2017; 2022).
-While approximate decision problems are still undecidable for general UMDPs over infinite horizons, Chatterjee et al.
-(2024) study a special case of UMDPs whose approximate decision problems are decidable.
-Burago et al.
-(1996) prove that computing the optimal policy for a POMDP over a finite horizon is NP-hard but showing that it is NP-hard for an UMDP.
-Wu & He (2020) study special cases of UMDPs over finite horizons whose optimal policies can be computed in polynomial time.
-UMDPs have also been used as approximations of POMDPs Hauskrecht (2000); Brechtel (2015); Lauri (2016) or simply discussed as a special case of POMDPs V (2009).
+The undecidability also holds for a restricted class of UMDPs [12, 13].
+While approximate decision problems are still undecidable for general UMDPs over infinite horizons, [29] study a special case of UMDPs whose approximate decision problems are decidable.
+[26] prove that computing the optimal policy for a POMDP over a finite horizon is NP-hard but showing that it is NP-hard for an UMDP.
+[139] study special cases of UMDPs over finite horizons whose optimal policies can be computed in polynomial time.
+UMDPs have also been used as approximations of POMDPs [54, 24, 79] or simply discussed as a special case of POMDPs [125].
 For a given POMDP, the corresponding UMDP can given a lower bound on the value function, while the corresponding (fully observable) MDP can give an upper bound on the value function.
@@
 Notice that an UMDP can allow more efficient optimization than the corresponding POMDP, since the UMDP does not need to deal with observations.
-For exemple, King (2018) studies an MCTS method for UMDPs.
-UMDPs have also been studied as a simple special case of POMDPs to study the effectiveness of planning methods in belief states to study the relative performance of different planning methods for POMDPs Littlefield (2020); Littlefield et al.
-(2018); Kimmel et al.
-(2019).
-UMDPs have also been simply discussed as a special case of POMDPs Boutilier et al.
-(1999); Csáji (2008); Verma & Rao (2005).
+For exemple, [73] studies an MCTS method for UMDPs.
+UMDPs have also been studied as a simple special case of POMDPs to study the effectiveness of planning methods in belief states to study the relative performance of different planning methods for POMDPs [81, 82, 72].
+UMDPs have also been simply discussed as a special case of POMDPs [22, 34, 126].
 UMDP also appears in the study of planning for multiple distributed agents to optimize a single objective under partial observability.
-Specifically, planning for a decentralized POMDP (DecPOMDP) Oliehoek & Amato (2016) can be reduced to planning for a (centralized) UMDP Oliehoek & Amato (2014); Roijers (2016); Roijers & Whiteson (2020), where the state in the UMDP is the pair of the state and the history of observations in the DecPOMDP, and the action in the UMDP is the decision rule that maps the history of observations into the actions in the POMDP.
+Specifically, planning for a decentralized POMDP (DecPOMDP) [93] can be reduced to planning for a (centralized) UMDP [92, 108, 109], where the state in the UMDP is the pair of the state and the history of observations in the DecPOMDP, and the action in the UMDP is the decision rule that maps the history of observations into the actions in the POMDP.
 In DecPOMDPs, the belief (distribution) over the pair of the state and the history of observations is the sufficient statistic, and planning can be performed in the space of the belief states.
-Even-Dar et al.
-(2007) consider an UMDP in the context of studying what values observations can provide in a POMDP.
+[41] consider an UMDP in the context of studying what values observations can provide in a POMDP.
 They introduce a parameter that ranges from 0 to 1.
 When the parameter is 0, the observation provides no information about the state (hence, the POMDP reduces to an UMDP); when the parameter is 1, the observation provides full information about the state (hence, the POMDP reduces to an MDP).
-The prior work also extends UMDPs to allow often costly actions that enable partial or full observations of the state Fox & Tennenholtz (2007); Kamar et al.
-(2009); Kamar (2010); Wang et al.
-(2025).
+The prior work also extends UMDPs to allow often costly actions that enable partial or full observations of the state [43, 68, 69, 130].